### v1

# Disentangled Representation Learning for Parametric Partial Differential Equations

Abstract
Neural operators (NOs) have demonstrated remarkable success in learning mappings between function spaces, serving as efficient approximators for the forward solutions of complex physical systems governed by partial differential equations (PDEs). However, while effective as black-box solvers, they offer limited insight into the underlying physical mechanism, due to the lack of interpretable representations of the physical parameters that drive the system. To tackle this challenge, we propose a new paradigm for learning disentangled representations from neural operator parameters, thereby effectively solving an inverse problem. Specifically, we introduce DisentangO, a novel hyper-neural operator architecture designed to unveil and disentangle the latent physical factors of variation embedded within the black-box neural operator parameters. At the core of DisentangO is a multi-task neural operator architecture that distills the varying parameters of the governing PDE through a task-wise adaptive layer, coupled with a hierarchical variational autoencoder that disentangles these variations into identifiable latent factors. By learning these disentangled representations, our model not only enhances physical interpretability but also enables more robust generalization across diverse physical systems. Empirical evaluations across supervised, semi-supervised, and unsupervised learning contexts show that DisentangO effectively extracts meaningful and interpretable latent features, bridging the divide between predictive performance and physical understanding in neural operator frameworks.
1 Introduction
Interpretability in machine learning (ML) refers to the ability to understand and explain how models make decisions (Rudin et al., 2022; Molnar, 2020). As ML systems grow more complex, especially with the use of deep learning and ensemble methods (Sagi and Rokach, 2018), the reasoning behind their predictions can become opaque. Interpretability addresses this challenge by making the models more transparent, enabling users to trust the outcomes, detect biases, and identify potential flaws. It is a critical factor in applying AI responsibly, especially in fields where accountability and fairness are essential (Cooper et al., 2022). In physics, where models endeavor to capture the governing laws and physical principles, understanding how a model arrives at its predictions is vital for verifying that it aligns with known scientific theories. Interpretable models also allow researchers to gain insights into the learned relationships, ensuring that the predictions are physically meaningful and consistent with real-world behaviors. This transparency is key for advancing scientific discovery, validating results, and enhancing trust in models used for complex physical systems.
Discovering interpretable representations of physical parameters in learning physical systems is challenging due to the intricate nature of real-world phenomena and the often implicit relationships between variables. In physics, parameters like force, energy, and velocity are governed by well-established laws, and extracting these parameters in a way that aligns with physical intuition requires models that go beyond mere pattern recognition. Traditional ML models may fit the data but fail to represent these parameters in a physically interpretable way. To address this, recent developments include the integration of physical constraints into the learning process (Raissi et al., 2019), such as embedding conservation laws (Liu et al., 2023, 2024a), symmetries (Mattheakis et al., 2019), or invariances (Cohen and Welling, 2016) directly into model architectures. Additionally, methods like symbolic regression (Biggio et al., 2021) and sparse modeling (Carroll et al., 2009) aim to discover simple, interpretable expressions that capture the underlying dynamics. However, balancing model complexity with interpretability remains a significant challenge, as overly complex models may obscure the true physical relationships, while oversimplified models risk losing critical details of the system’s behavior.
In this work, we introduce DisentangO, a novel variational hierarchical neural operator architecture to disentangle physical factors of variation from black-box neural operator parameters for solving parametric PDEs. Neural operators (NOs) (Li et al., 2020a, c) learn mappings between infinite-dimensional function spaces in the form of integral operators, making them powerful tools for discovering continuum physical laws by manifesting the mappings between spatial and/or spatiotemporal data; see You et al. (2022a); Liu et al. (2024a, b, 2023); Ong et al. (2022); Cao (2021); Lu et al. (2019, 2021); Goswami et al. (2022); Gupta et al. (2021) and references therein. However, most NOs play the role of efficient forward surrogates for the underlying physical system under a supervised learning setting. As a result, they act as black-box universal approximators for a single physical system, and are lack of interpretability with respect to the underlying physical laws. In contrast, the key innovation of DisentangO lies in its use of a hierarchical architecture that distills the varying physical parameters of the governing PDE from multiple physical systems through an unsupervised learning setting. The distilled variations are further disentangled into distinct physical factors to enhance physical understanding and promote robust generalization. Consequently, DisentangO effectively extracts meaningful and interpretable physical features, thereby simultaneously solving both the forward and inverse problems. Our key contributions are:
  • We bridge the divide between predictive accuracy and physical interpretability, and introduce a new paradigm that simultaneously performs physics modeling (i.e., as a forward PDE solver) and governing physical mechanism discovery (i.e., as an inverse PDE solver).
  • We propose a novel variational hierarchical neural operator architecture, which we coin DisentangO. DisentangO extracts the key physical factors of variation from black-box neural operator parameters of multiple physical systems. These factors are then disentangled into distinct latent factors that enhance physical interpretation and robust generalization.
  • We provide theoretical analysis on the component-wise identifiability of the true generative factors in physics modeling: by learning from multiple physical systems, the variability of hidden physical states in these systems promotes identifiability.
  • We explore the practical utility of disentanglement rather than pursuing it as an end goal, and perform experiments across a broad range of settings including supervised, semi-supervised, and unsupervised learning. Our results show that DisentangO effectively extracts meaningful and interpretable physical features for enhanced interpretability.
2 Background and related work
Neural operators. Learning complex physical systems directly from data is crucial in many scientific and engineering applications (Ghaboussi et al., 1998; Liu et al., 2024c; Ghaboussi et al., 1991; Carleo et al., 2019; Karniadakis et al., 2021; Zhang et al., 2018; Cai et al., 2022; Pfau et al., 2020; He et al., 2021; Jafarzadeh et al., 2024). Often, the governing laws of these systems are unknown and must be inferred from data through learning models that are both resolution-invariant to ensure consistent performance across various levels of discretizations and interpretable to domain experts. Neural operators (NOs) address the first aim by learning mappings between infinite-dimensional function spaces (Li et al., 2020a, b, c; You et al., 2022a; Ong et al., 2022; Cao, 2021; Lu et al., 2019, 2021; Goswami et al., 2022; Gupta et al., 2021). By capturing the relationship between spatial and spatio-temporal data, NOs offer a promising approach to discovering continuum physical surrogates, enabling more accurate and consistent predictions of the underlying responses of complex systems. Nevertheless, NOs lack an interpretable representation of the physical states driving the system, and as a result offer limited insight into the underlying physical mechanisms.
Forward and inverse PDE learning. Existing NOs focus on providing efficient surrogates for the underlying physical systems as forward PDE solvers. They are often employed as black-box universal approximators but lack interpretability of the underlying physical principles. On the other hand, several deep learning methods have been developed as inverse PDE solvers (Fan and Ying, 2023; Molinaro et al., 2023; Jiang et al., 2022; Chen et al., 2023), aiming to reconstruct the PDE parameters from solution data. However, the inverse problem is typically more challenging due to its ill-posedness. To tackle this pathology, many NOs incorporate prior information via governing PDEs (Yang et al., 2021; Li et al., 2021), regularizers (Dittmer et al., 2020; Obmann et al., 2020; Ding et al., 2022; Chen et al., 2023), or additional operator structures (Uhlmann, 2009; Lai et al., 2019; Yilmaz, 2001). However, all these work rely on prior information on the operator form, which is not realistic in many real-world problems where neither the parameter field nor the model form are available. Worth noting is a recent work that leverages the attention mechanism to build a nonlinear kernel map in NO (Yu et al., 2024), with which a data-dependent kernel can be constructed to characterize the inverse mapping from data to the hidden governing mechanism. While effective, this approach does not disentangle the learned kernel or extract physically interpretable parameters. To the best of our knowledge, our proposed DisentangO is the first that tackles both forward (i.e., physics prediction) and inverse (i.e., physics discovery) PDE learning problems, while simultaneously inferring the governing distinct physical parameters from the learned NO.
Disentangled representation learning. Disentangled representation learning seeks to separate data into distinct, interpretable factors, where each factor captures an independent aspect of variation in the input. This area has critical implications for transfer learning, generative modeling, and fairness in AI, as it allows models to generalize across tasks by leveraging isolated features, making them more interpretable and modular. Disentangled representation learning has seen significant advances, primarily driven by models like $\beta$-VAE (Higgins et al., 2017), FactorVAE (Kim and Mnih, 2018), and InfoGAN (Chen et al., 2016), who encourage disentanglement by controlling the trade-off between reconstruction accuracy and latent regularization, or introduce mutual information constraints to encourage the discovery of interpretable latent factors. Early work in disentanglement largely focus on supervised or semi-supervised approaches (Ridgeway and Mozer, 2018; Shu et al., 2019; Mathieu et al., 2019), requiring labeled data for factors of variation. However, more recent advancements have concentrated on unsupervised learning methods (Duan et al., 2019), where disentanglement is achieved without explicit labels. Despite these efforts, fully unsupervised disentanglement remains challenging (Locatello et al., 2019), with ongoing research focusing on improving metrics, robustness, and applicability to more complex and high-dimensional data. In this context, although a large body of research has been done in disentanglement, its applications remain largely confined to fields like computer vision and robotics, where the latent dimensions typically correspond to human-interpretable (often visually recognizable) features, with very limited exploration in areas such as physical system learning. Moreover, existing work focuses on disentangling from data, whereas our work is the first attempt that explores disentanglement from black-box neural network parameters.
3 DisentangO: disentangled representation learning from neural operator parameters
We consider a series of complex systems with different hidden physical parameters:
$\mathcal{K}_{\bm{b}}[\bm{u}](\bm{x})=\bm{f}(\bm{x})\text{ ,}\quad\bm{x}\in{\Omega}\text{ .}$ (3.1)
Here, ${\Omega}\subset\mathbb{R}^{s}$ is the domain of interest, $\bm{f}(\bm{x})$ is a function representing the loading on ${\Omega}$, and $\bm{u}(\bm{x})$ is the corresponding solution of this system. $\mathcal{K}_{\bm{b}}$ represents the unknown governing law, e.g., balance laws, which is determined by the (possibly unknown and high-dimensional) parameter field $\bm{b}$. For instance, in a material modeling problem, $\mathcal{K}_{\bm{b}}$ often stands for the constitutive law and $\bm{b}$ can be a vector ($\bm{b}\in\mathbb{R}^{d_{b}}$) representing the homogenized material parameter field or a vector-valued function ($\bm{b}\in L^{\infty}({\Omega};\mathbb{R}^{d_{b}})$) representing the heterogeneous material properties. Both scenarios are considered in our empirical experiments in Section 4.
Many physical modeling tasks can be formulated as either a forward or an inverse PDE-solving problem. In a forward problem setting, one aims to find the PDE solution when given PDE information, including coefficient functions, boundary conditions, initial conditions, and loading sources. That means, given the governing operators $\mathcal{K}$, the parameter (field) $\bm{b}$, and loading field $\bm{f}(\bm{x})$ in equation 3.1, the goal is to solve for the corresponding solution field $\bm{u}(\bm{x})$, through classical PDE solvers (Brenner and Scott, 2007) or data-driven approaches (Lu et al., 2019; Li et al., 2020c; You et al., 2022b). As a result, a forward map is constructed:
[…]
Here, $\bm{b}$ and $\bm{f}$ are input vectors/functions, and $\bm{u}$ is the output function.
Conversely, solving an inverse PDE problem involves reconstructing the underlying full or partial PDE information from PDE solutions, where one seeks to construct an inverse map:
[…]
Unfortunately, solving an inverse problem is typically more challenging due to the ill-posed nature of the PDE model. In general, a small number of function pairs $(\bm{u},\bm{f})$ from a single system does not suffice in inferring the underlying parameter field $\bm{b}$, making the inverse problem generally non-identifiable (Molinaro et al., 2023). Herein, we propose a novel neural architecture to alleviate the curse of ill-posedness without requiring prior knowledge. The key ingredients are: 1) the construction of a multi-task NO architecture, which solves both the forward (physics prediction) and inverse (physics discovery) problems simultaneously from multiple systems, and 2) a generative model to disentangle the features from NO parameters that contain critical information of $\bm{b}$.
3.1 Variational hierarchical neural operator as a multi-task solver
Figure 1: Illustration of DisentangO’s architecture.
Formally, we consider $S$ training datasets from different systems, each of which contains $n_{train}$ function pairs $(\bm{u},\bm{f})\in\mathcal{U}\times\mathcal{F}$:
$\mathcal{D}_{\rm tr}=\{\{(\bm{u}^{\eta}_{i}(\bm{x}),\bm{f}^{\eta}_{i}(\bm{x}))\}_{i=1}^{n_{train}}\}_{\eta=1}^{S}\text{ .}$ (3.4)
The index $\eta$ corresponds to different hidden parameter field/set $\bm{b}^{\eta}\in\mathcal{B}$. Here, $\mathcal{U}\subset L^{2}({\Omega};\mathbb{R}^{d_{u}})$, $\mathcal{F}\subset L^{2}({\Omega};\mathbb{R}^{d_{f}})$ represent the (infinite-dimensional) Banach spaces of the solution function $\bm{u}$ and loading function $\bm{f}$, respectively, and $\mathcal{B}$ is a finite-dimensional manifold where the hidden parameter $\bm{b}$ takes values on. We further assume that the data satisfies the following generating process:
$\bm{b}|\bm{z}=m(\bm{z})\text{ ,}\;\bm{z}\sim\mathbb{P}_{z}\text{ ,}$ (3.5)
where $\bm{z}\in\mathcal{Z}\subset\mathbb{R}^{d_{z}}$ is the embedding of $\bm{b}$ in the latent space and $m:\mathcal{Z}\rightarrow\mathcal{B}$ is a one-to-one function. Then, we assume for each $\bm{b}^{\eta}$ that there exists an underlying parametric PDE system of equation 3.1. Hence, each function pair $(\bm{u}_{i}^{\eta},\bm{f}^{\eta}_{i})$ satisfies:
$\bm{u}|\bm{f}=\mathcal{K}_{\bm{b}}^{-1}[\bm{f}]\text{ ,}\;\bm{f}\sim\mathbb{P}_{f}\text{ ,}$ (3.6)
where $\mathcal{K}_{\bm{b}}^{-1}$ is the ground-truth solution operator of the parametric PDE system of equation 3.1 with a given parameter $\bm{b}$.
Our first objective is to construct a surrogate operator of $\mathcal{K}_{\bm{b}}^{-1}:\mathcal{F}\rightarrow\mathcal{U}$, as the forward PDE solver for each system:
$\mathcal{G}[\bm{f}^{\eta}_{i};\theta^{\eta}](\bm{x})\approx\mathcal{K}_{\bm{b}^{\eta}}^{-1}[\bm{f}_{i}^{\eta}](\bm{x})=\bm{u}_{i}^{\eta}(\bm{x})\text{ ,}$ (3.7)
where $\theta^{\eta}$ denotes the corresponding trainable parameter set. Then, as the second objective, we construct an inverse map by discovering the hidden parameter field/set from $\theta^{\eta}$:
$\mathcal{H}(\theta^{\eta};\Theta_{H})\approx\tilde{\bm{z}}^{\eta}\text{ .}$ (3.8)
Here, $\Theta_{H}$ is the trainable parameters of $\mathcal{H}$, and $\tilde{\bm{z}}^{\eta}\in\mathbb{R}^{d_{z}}$ denotes the learned latent variables, which can be transformed to the ground-truth latent variable $\bm{z}^{\eta}$ via an invertible function $h$. The goal of disentanglement is not to discover $\bm{b}^{\eta}$, since some features may not be identifiable. For instance, in a Dirichlet boundary condition problem, $\bm{u}^{\eta}_{i}(\bm{x})=0$ for all $\bm{x}\in\partial\Omega$, making $\bm{b}^{\eta}$ not learnable on $\partial\Omega$. Hence, a more realistic goal is to discover the underlying mechanism of $\bm{b}$ in the space of identifiability, i.e., $\bm{z}$, or its equivalent transformation $\tilde{\bm{z}}$.
Note that in the first objective of equation 3.7, one is interested in finding the solution operator $\mathcal{G}^{\eta}$ under a supervised setting, in the form of function-to-function mappings, for all (hidden) parameters $\bm{b}^{\eta}$ in the range of interest; while in the second objective of equation 3.8, the goal is to construct a vector-to-vector mapping from $\theta^{\eta}$ to $\tilde{\bm{z}}^{\eta}$, which is unsupervised due to the hidden physics nature. We propose to employ variational autoencoder (VAE) (Kingma and Welling, 2013) as the representation learning approach for the second objective, and a meta-learned NO architecture as a universal solution operator for the first objective. The key is to pair these architectures as a hierarchical neural operator, which simultaneously solves both forward and inverse modeling problems.
We now define the learning architecture. The overall objective is to maximize the log data likelihood:
$\max E(\log(p(\theta,\mathcal{D})))=\max\left[E(\log(p(\bm{u}|\bm{f},\theta)+\log(p(\theta|\bm{f}))+\log(p(\bm{f}))\right].$
[…]
However, the second term in this formulation is generally intractable. We use a variational posterior $q(\theta|\bm{z})$ to approximate the actual posterior $p(\theta|\bm{z})$ and maximize the evidence lower bound (ELBO):
$L_{ELBO}=\dfrac{1}{S}\sum_{\eta=1}^{S}\left[E_{q(\bm{z}^{\eta}|\theta^{\eta})}\log p(\theta^{\eta}|\bm{z}^{\eta})-D_{KL}\left(q(\bm{z}^{\eta}|\theta^{\eta})||p(\bm{z}^{\eta})\right)\right].$
Putting everything together, we obtain the loss functional:
$L_{loss}=\dfrac{1}{S}\sum_{\eta=1}^{S}\left[-E(\log(p(\bm{u}|\bm{f},\theta^{\eta})))-E_{q(\bm{z}^{\eta}|\theta^{\eta})}\log p(\theta^{\eta}|\bm{z}^{\eta})+D_{KL}\left(q(\bm{z}^{\eta}|\theta^{\eta})||p(\bm{z}^{\eta})\right)\right].$ (3.10)
The above formulation naturally lends itself to a hierarchical variational autoencoder (HVAE) architecture (Vahdat and Kautz, 2020). The encoder aims to obtain the posterior $q_{\mu_{z},\Sigma_{z}}(\bm{z}^{\eta}|\theta^{\eta})$, which estimates the inverse map $\mathcal{H}$:
$\hat{\bm{z}}\sim q_{\mu_{z},\Sigma_{z}}(\hat{\bm{z}}^{\eta}|\theta^{\eta})\text{ .}$ (3.11)
Then, the first decoder, which processes the estimated latent variable $\hat{\bm{z}}$ and reconstructs the corresponding NO parameter $\hat{\theta}$:
$\hat{\theta}=\hat{g}(\hat{\bm{z}})\text{ .}$ (3.12)
Lastly, the second decoder reconstructs the forward map $\mathcal{G}$, by further taking the estimated $\hat{\theta}$ and the loading function $\bm{f}$ and estimating the output function $\bm{u}$:
$\hat{\bm{u}}=\hat{\mathcal{G}}[\bm{f};\hat{\theta}]\text{ .}$ (3.13)
For this purpose, we employ the implicit Fourier neural operator (IFNO)***We also point out that the proposed strategy is generic and hence also applicable to other neural operators., which is a provably universal approximator for PDE solution operators with a relatively small number of trainable parameters (You et al., 2022b). To provide a universal $\hat{\mathcal{G}}$ for different domains, a multi-task learning strategy is adopted following Zhang et al. (2023), which applies task-wise adaptation only to the first layer of NOs. Specifically, an $L$-layer multi-task IFNO has the following form:
${\mathcal{G}}[\bm{f};{\theta}^{\eta}](\bm{x})={\mathcal{G}}[\bm{f};{\theta}_{P}^{\eta},\theta_{J},\theta_{Q}](\bm{x}):=\mathcal{Q}_{\theta_{Q}}\circ(\mathcal{J}_{\theta_{J}})^{L}\circ\mathcal{P}_{{\theta}^{\eta}_{P}}[\bm{f}](\bm{x})\text{ ,}$ (3.14)
where $\mathcal{P}$, $\mathcal{Q}$ are shallow-layer MLPs that map a low-dimensional vector into a high-dimensional vector and vice versa, parameterized by $\theta_{Q}$ and ${\theta}^{\eta}_{P}$, respectively. Each intermediate layer, $\mathcal{J}$, is constructed as a mimetic of a fixed-point iteration step and parameterized by $\theta_{J}$. Supported by the universal approximator analysis in Zhang et al. (2023), different PDEs share the common iterative ($\theta_{J}$) and projection ($\theta_{Q}$) parameters, with all the information about parameter $\bm{b}$ embedded in the task-wise lifting parameters $\theta^{\eta}_{P}$. Hence, the encoder and first decoder structure only need to encode and reconstruct $\theta^{\eta}_{P}$, since all other NO parameters are invariant to the change of $\bm{z}$. This construction substantially reduces the degree of freedoms in $\theta^{\eta}$, making the invertibility assumption in the next section feasible. With a slight abuse of notation, in the following contents we use $\theta^{\eta}$ to denote the task-wise parameter, i.e., the lifting layer parameter $\theta^{\eta}_{P}$, and estimate $\bm{z}^{\eta}$ from it.
3.2 Disentangling the underlying mechanism
The proposed DisentangO aims to identify and disentangle the components of the latent representation $\bm{z}$, which serves as an inverse PDE solver. However, to capture the true physical mechanism, a natural question arises: is it really possible to identify the latent variables of interest (i.e., $\bm{z}$) with only observational data $\mathcal{D}_{\rm tr}$? We now show that, by learning the model $(p_{\hat{\bm{z}}},\hat{g},\hat{\mathcal{H}},\hat{\mathcal{G}})$ that matches the true marginal data distribution in all domains, we can indeed achieve this identifiability for the generating process proposed in equation 3.5 and equation 3.6.
First, we show that the latent variable $\bm{z}$ can be identified up to an invertible component-wise transformation: for the true latent variable $\bm{z}\in\mathbb{R}^{d_{z}}$, there exists an invertible function $h:\mathbb{R}^{d_{z}}\rightarrow\mathbb{R}^{d_{z}}$, such that $\hat{\bm{z}}=h(\bm{z})$.
Theorem 1.
We follow the data-generating process in equation 3.5 and equation 3.6, and make the following assumptions:
  • (Density Smoothness and Positivity): $p_{\theta}$ and $p_{\bm{z}}$ are both smooth and positive.
  • (Invertibility): The task-wise parameter $\theta$ can be generated by $\bm{z}$ through an invertible and smooth function $\mathcal{H}^{-1}$. Moreover, for each given $\bm{f}$, we denote $\mathcal{G}_{f}(\theta):=\mathcal{G}[\bm{f};\theta](\bm{x})$ as the operator mapping from $\mathbb{R}^{d_{\theta}}$ to $\mathcal{U}$. $\mathcal{G}_{f}$ is also one-to-one with respect to $\theta$.
Then, by learning $(p_{\hat{\bm{z}}},\hat{g},\hat{\mathcal{H}},\hat{\mathcal{G}})$ to achieve:
$p_{\hat{\bm{u}}|\bm{f}}=p_{\bm{u}|\bm{f}}\text{ ,}$ (3.15)
where $\bm{u}$ and $\hat{\bm{u}}$ are generated from the true process and the estimated model, respectively, $\bm{z}$ is identifiable up to an invertible function $h$.
Proof: Please see Appendix A.
Additionally, with additional assumptions on conditional independence and datum variability, we can further obtain the following theoretical results on component-wise identifiability: for each true component $z_{i}$, there exists a corresponding estimated component $\hat{z}_{j}$ and an invertible function $h_{i}:\mathbb{R}\rightarrow\mathbb{R}$, such that $z_{i}=h_{i}(\hat{z}_{j})$.
Theorem 2.
We follow the data-generating process in equation 3.5 and equation 3.6, and assume that the Density Smoothness and Positivity condition as well as the Invertibility condition in Theorem 1 are both satisfied, together with the following additional assumptions:
  • (Conditional Independence): Conditioned on $\bm{b}$, each component of $\bm{z}$ is independent of each other: $\log p_{\bm{z}|\bm{b}}(\bm{z}|\bm{b})=\sum_{i=1}^{d_{z}}\log p_{z_{i}|\bm{b}}(z_{i}|\bm{b})$.
  • (Linear independence) There exists $12d_{z}+$ 1 values of $\bm{b}$, such that the $2d_{z}$ vectors $w(\bm{z},\bm{b}^{j})-w(\bm{z},\bm{b}^{0})$ with $j=1,\cdots,2d_{z}$ are linearly independent. Here,
    $w(\bm{z},\bm{b}^{j}):=\left(\dfrac{\partial q_{1}(z_{1},\bm{b}^{j})}{\partial z_{1}},\cdots,\dfrac{\partial q_{d_{z}}(z_{d_{z}},\bm{b}^{j})}{\partial z_{d_{z}}},\dfrac{\partial^{2}q_{1}(z_{1},\bm{b}^{j})}{\partial z_{1}^{2}},\cdots,\dfrac{\partial^{2}q_{d_{z}}(z_{d_{z}},\bm{b}^{j})}{\partial z_{d_{z}}^{2}}\right),$ (3.16)
    with $q_{i}(z_{i},\bm{b}):=\log p_{z_{i}|\bm{b}}$.
Then, by learning $(p_{\hat{\bm{z}}},\hat{g},\hat{\mathcal{H}},\hat{\mathcal{G}})$ to achieve equation 3.15, $\bm{z}$ is component-wise identifiable.
Proof: Please see Appendix A.
Intuitively, the linear independence condition in Theorem 2 indicates sufficient variability between physical systems (tasks). This is a common assumption in the nonlinear ICA literature for domain adaptation (Hyvarinen et al., 2019; Khemakhem et al., 2020; Kong et al., 2023). To our best knowledge, this is the first time the component-wise identifiability results are discussed in the context of multi-task neural operator learning.
3.3 A generic algorithm for different supervision scenarios
In this section, we discuss the practical utility of DisentangO under three scenarios:
  • (SC1: Unsupervised) No value or label is given for $\bm{b}^{\eta}$ for each task.
  • (SC2: Semi-supervised) The value of $\bm{b}^{\eta}$ is not given, but a label $c(\bm{b}^{\eta})$ (e.g., a classification of $\bm{b}^{\eta}$) is given.
  • (SC3: Supervised) The value of $\bm{b}^{\eta}$ is given.
The architecture is shown in Figure 1, with a pseudo algorithm provided in Algorithm 1 of Appendix.
To obtain the posterior $q$ in equation 3.11, we assume each latent variable satisfies a Gaussian distribution of distinct means and diagonal covariance, i.e., $q_{\mu_{z},\Sigma_{z}}(\hat{\bm{z}}|\theta):=\mathcal{N}(\mu_{z}(\theta_{P}),\Sigma^{2}_{z}(\theta_{P}))$, then estimate its mean and covariance using an MLP. As a result, the KL-divergence term in the ELBO admits a closed form:
[…]
In the supervised setting, we take $\bm{z}=\bm{b}$, and set $\mu_{z}(\theta_{P}^{\eta})=\bm{b}^{\eta}$, then:
[…]
For the first decoder, the likelihood $p(\theta^{\eta}_{P}|\bm{z}^{\eta})$ is a factorized Gaussian with mean $\mu_{\theta}$ and covariance $\Sigma_{\theta}$, computed from another MLP. By taking as input Monte Carlo samples once for each $\bm{z}$, the reconstruction accuracy term can be approximated as:
$E_{q(\bm{z}|\theta_{P})}\log p(\theta_{P}|\bm{z})\approx-\sum_{i=1}^{d_{\theta}}\left(\dfrac{((\theta_{P})_{i}-(\mu_{\theta})_{i})^{2}}{2(\Sigma_{\theta})^{2}_{i}}+\log(\Sigma_{\theta})_{i}+c\right)\text{ ,}$
where ${d_{\theta}}$ is the dimension of $\theta_{P}$, and $c$ is a constant.
For the second decoder, we parameterize it as a multi-task NO in equation 3.14, and model the discrepancy between $\hat{\mathcal{G}}_{[\hat{\theta}^{\eta}_{P},\theta_{I},\theta_{Q}]}[\bm{g}^{\eta}_{j}](\bm{x}_{k})$ and the ground truth $\bm{u}(\bm{x}_{k})$ as an additive independent unbiased Gaussian random noise $\epsilon$, with
$\hat{\mathcal{G}}_{[\hat{\theta}^{\eta}_{P},\theta_{I},\theta_{Q}]}[\bm{g}^{\eta}_{j}](\bm{x}_{k}))=\bm{u}_{j}^{\eta}(\bm{x}_{k})+\epsilon_{\eta,j,k}\text{ ,}\quad\epsilon_{\eta,j,k}\sim\mathcal{N}(0,\varpi^{2})\text{ ,}$
where $\varpi$ is a tunable hyperparameter. Then, with a uniform spatial discretization of size $\Delta x$ in a domain ${\Omega}\subset\mathbb{R}^{d_{u}}$, the log likelihood after eliminating the constant terms can be written as:
$\dfrac{1}{2\varpi^{2}\Delta x^{2d_{u}}}\sum_{\eta=1}^{S}\sum_{j=1}^{n_{train}}{\left|\left|\hat{\mathcal{G}}_{[\hat{\theta}^{\eta}_{P},\theta_{I},\theta_{Q}]}[\bm{g}^{\eta}_{j}](\bm{x}_{k})-\bm{u}_{j}^{\eta}(\bm{x}_{k})\right|\right|}_{L^{2}({\Omega})}^{2}.$
In our empirical tests in Section 4, for the unsupervised and semi-supervised scenarios we choose $q$ as a standard Gaussian distribution following the independence assumption of Theorem 2. Additionally, to avoid overparameterization, we take $\Sigma_{\theta}=\sigma^{2}_{\theta}\mathbf{I}$, with $\sigma_{\theta}$ being a tunable hyperparameter. In this case, the total objective can be further simplified as:
$L_{loss}=$ $\dfrac{1}{S}\sum_{\eta=1}^{S}\left(\beta_{d}\sum_{j=1}^{n_{train}}{\left|\left|\hat{\mathcal{G}}_{[\hat{\theta}^{\eta}_{P},\theta_{I},\theta_{Q}]}[\bm{g}^{\eta}_{j}]-\bm{u}^{\eta}_{j}\right|\right|}^{2}_{L^{2}({\Omega})}+{\left|\left|\hat{\theta}^{\eta}_{P}-\mu^{\eta}_{\theta}\right|\right|}^{2}+\beta_{KL}{\left|\left|\mu^{\eta}_{\bm{z}}\right|\right|}^{2}\right)\text{ ,}$
for the unsupervised scenario, and
$L_{loss}=$ $\dfrac{1}{S}\sum_{\eta=1}^{S}\left(\beta_{d}\sum_{j=1}^{n_{train}}{\left|\left|\hat{\mathcal{G}}_{[\hat{\theta}^{\eta}_{P},\theta_{I},\theta_{Q}]}[\bm{g}^{\eta}_{j}]-\bm{u}^{\eta}_{j}\right|\right|}^{2}_{L^{2}({\Omega})}+{\left|\left|\hat{\theta}^{\eta}_{P}-\mu^{\eta}_{\theta}\right|\right|}^{2}+\beta_{KL}{\left|\left|\mu^{\eta}_{\bm{z}}-\bm{b}^{\eta}\right|\right|}^{2}\right)\text{ ,}$
for the fully supervised scenario. Here, $\beta_{d}:=\frac{\sigma^{2}_{\theta}}{\varpi^{2}\Delta x^{2d_{u}}}$ and $\beta_{KL}:=\sigma^{2}_{\theta}$ are treated as tunable hyperparameters. In the semi-supervised scenario, we incorporate partial supervision by equipping the above loss with a constraint following Locatello et al. (2020):
$L_{loss}=\dfrac{1}{S}\sum_{\eta=1}^{S}$ $\left(\beta_{d}\sum_{j=1}^{n_{train}}{\left|\left|\hat{\mathcal{G}}_{[\hat{\theta}^{\eta}_{P},\theta_{I},\theta_{Q}]}[\bm{g}^{\eta}_{j}]-\bm{u}^{\eta}_{j}\right|\right|}^{2}_{L^{2}({\Omega})}+{\left|\left|\hat{\theta}^{\eta}_{P}-\mu^{\eta}_{\theta}\right|\right|}^{2}\right.+\left.\beta_{KL}{\left|\left|\mu^{\eta}_{\bm{z}}\right|\right|}^{2}+\beta_{cls}L_{c}(\bm{z}^{\eta},c(\beta^{\eta}))\right)$
where $\beta_{cls}$ is a tunable hyperparameter, and $L_{c}(\bm{z}^{\eta},c(\beta^{\eta}))$ is the corresponding loss term depending on the provided information. For instance, when taking $c$ as the classification label of $\beta^{\eta}$, $L_{c}$ can be taken as the cross-entropy loss.
We note that the coefficient $\beta_{KL}$ is closely connected to the adjustable hyperparameter in $\beta-$VAE (Higgins et al., 2017). Taking a larger $\beta_{KL}$ encourages disentanglement (Locatello et al., 2019), while it also adds additional constraint on the implicit capacity of the latent bottleneck that results in information loss (Burgess et al., 2018). On the other hand, the data reconstruction term forces the latent factors to contribute to the “reconstruction” of the more complicated solution field in a global perspective, and hence increasing $\beta_{d}$ is anticipated to alleviate information loss. In our empirical studies, we demonstrate the interplay between these two tunable hyperparameters.
4 Experiments
We assess the performance of DisentangO across a variety of physics modeling and discovery datasets. In the following, we denote a DisentangO model with a latent dimension of $n$ as DisentangO-$n$. Our evaluation focuses on the following key aspects. Firstly, we demonstrate the capability of DisentangO in forward PDE learning, and compare its performance with a number of relevant baselines. In particular, we select three NO-related baselines, i.e., FNO (Li et al., 2020c), NIO (Molinaro et al., 2023) and MetaNO (Zhang et al., 2023), as well as two non-NO baselines, i.e., VAE (Kingma and Welling, 2013) and $\beta$-VAE (Higgins et al., 2017). Secondly, we showcase the merits of DisentangO in inverse modeling for interpretable physics discovery. Thirdly, we perform parametric studies on the associated disentanglement hyperparameters. Lastly, we provide interpretations of the disentangled parameters and point out future directions. Further details are provided in Appendix C.
4.1 Supervised forward and inverse PDE learning
Table 1: Test errors and the number of trainable parameters. Bold number highlights the best method.
Models #param Test errors
data latent supervision
DisentangO 697,383 1.65% 4.63%
MetaNO 295,842 1.59% -
NIO 709,260 - 15.16%
FNO 698,284 2.45% 14.55%
VAE 697,672 49.97% 16.34%
$\beta$-VAE 697,672 51.16% 16.47%
We start by investigating DisentangO’s capability in solving both forward and inverse PDE problems in a fully supervised setting. Specifically, we consider the constitutive modeling of anisotropic fiber-reinforced hyperelastic materials governed by the Holzapfel–Gasser–Ogden (HGO) model, where the data-generating process is controlled by sampling the governing material parameter set $\{E,\nu,k_{1},k_{2},\alpha\}$ and the latent factors can be learned consequently in a supervised fashion. In this setting, the model takes as input the padded traction field and learns to predict the displacement.
As the number of latent factors is fixed for supervised learning in this example, we defer our ablation study to the second experiment of semi-supervised learning. We report in Table 1 our experimental results from DisentangO as well as its comparison with the selected baselines. Note that only FNO, VAE and $\beta$-VAE are able to handle both forward and inverse problems, whereas MetaNO and NIO can only solve the forward problem and the inverse problem, respectively. With this caveat in mind, MetaNO achieves the best performance in forward PDE learning, with DisentangO performing comparably well and beating the third best model by 32.7% in accuracy. In terms of inverse modeling with full latent supervision, DisentangO is the only method that can hold the error below 5%, outperforming the second best method by 68.2%.
4.2 Semi-supervised mechanical MNIST benchmark
We next explore in the context of semi-supervised learning and apply DisentangO and relevant baselines to the benchmark Mechanical MNIST (MMNIST) dataset (Lejeune, 2020). MMNIST comprises 70,000 heterogeneous material specimens undergoing large deformation, each governed by a material model of the Neo-Hookean type with a varying modulus converted from the MNIST bitmap images. In our experiment, we take 500 images (with a 420/40/40 split for training/validation/test) and generate 200 loading/response data pairs per sample on a $29\times 29$ structured grid, simulating uniaxial extension, shear, equibiaxial extension, and confined compression load scenarios. Since only partial knowledge is available for each image (i.e., the corresponding digit), we apply semi-supervised learning to the latent factors to classify the digits.
Ablation study. Firstly, we investigate DisentangO’s predictability in forward PDE learning by comparing its performance to MetaNO (i.e., the base meta-learnt NO model without the hierarchical structure). As seen in Table 2, MetaNO achieves a forward prediction error of 2.68%, which serves as the optimal bound for DisentangO. As we increase the latent dimension in DisentangO from 2 to 15, the prediction error drops from 11.49% to 5.48% and converges to the optimal bound. Next, we study the role of the data loss term in disentanglement by gradually varying $\beta_{d}$ from $\beta_{d}=1$ to $\beta_{d}=1000$. In Table 2 we observe a consistent improvement in accuracy with the increase in $\beta_{d}$, where the boost in accuracy becomes marginal or slightly deteriorates beyond $\beta_{d}=100$. We thus choose $\beta_{d}=100$ as the best DisentangO model in this case. Besides offering an increased accuracy in forward modeling, the data loss term also enhances disentanglement as discussed in Section 3.2. This is evidenced by the illustration in Figure 2, where the unsupervised mutual information (MI) score that measures the amount of MI across latent factors consistently decreases in both DisentangO-2 and DisentangO-15 as we increase $\beta_{d}$ from $\beta_{d}=1$ (solid lines) to $\beta_{d}=100$ (dashed lines). On the contrary, the classification term poses a negative effect on disentanglement, as indicated by the increase in $\beta_{\text{cls}}$ leading to an increase in MI scores. This is reasonable because the classifier linearly combines all the latent factors to make a classification. The more accurate the classification, the stronger the correlation across latent factors. We then move on to study the effect of semi-supervised learning by comparing the model’s performance with and without latent partial supervision. While the models without semi-supervision is slightly more accurate in that the forward prediction accuracy with $\beta_{d}=1$ reaches 12.60% and 6.16% in DisentangO-2 and DisentangO-15, respectively, the latent scatterplot in Figure 3 reveals the inability of the DisentangO model without latent semi-supervision to acquire the partial knowledge of the embedded digits from data. In contrast, although the accuracy of DisentangO with latent semi-supervision slightly deteriorates due to the additional regularization effect from the classification loss term, it is able to correctly recognize the embedded digits and leverage this partial knowledge in disentangling meaningful latent factors.
Table 2: Test errors and the number of trainable parameters in MMNIST benchmark. Models are abbreviated (0=DisentangO-2, 1=DisentangO-5, 2=DisentangO-10, 3=DisentangO-15, 4=VAE, 5=$\beta$-VAE, 6=MetaNO). Bold number highlights the best method that can handle both forward and inverse settings.
Models 0 1 2 3 4 5 6
#param (M) 0.66 0.97 1.49 2.02 2.02 2.02 0.35
Error $\beta_{d}=1$ 12.82% 9.56% 7.36% 6.29% 16.34% 17.13% 2.68%
$\beta_{d}=10$ 11.51% 9.16% 6.62% 5.95% - - -
$\beta_{d}=100$ 11.49% 8.43% 6.65% 5.48% - - -
$\beta_{d}=1000$ 11.62% 8.22% 6.50% 5.80% - - -
Figure 2: MMNIST unsupervised scores against $\beta_{d}$ with DisentangO-2 (left) and DisentangO-15 (right). By comparing $\beta_{d}=1$ (solid lines) with $\beta_{d}=100$ (dashed lines), increasing $\beta_{d}$ forces the latent factors to maximize the contained information and in turn decreases MI, thus encouraging disentanglement.
Figure 3: MMNIST scatterplot with DisentangO-2 and $\beta_{d}=1$: left: $(\beta_{kl}=1,\beta_{cls}=100)$ (data test error 18.81%), middle: $(\beta_{kl}=10,\beta_{cls}=10)$ (data test error 16.94%), right: fully unsupervised DisentangO-2 without classification loss $(\beta_{kl}=100)$ (data test error 12.65%).
Comparison with additional baselines. We compare the performance of DisentangO with relevant additional baselines, i.e., VAE and $\beta$-VAE. We abandon the three NO-related baselines in this and the following experiments because they are black-box approximators and do not possess any mechanism to extract meaningful information without supervision. In this context, even the least accurate DisentangO-2 model (i.e., with $\beta_{d}=1$) with significantly fewer parameters outperforms the two selected baselines in Table 2 by 21.5% and 25.2%, respectively, with the best performing DisentangO-15 model beating the two baselines by 66.5% and 68.0% in accuracy, respectively. We do not proceed to compare the physics discovery capability between them as the baseline models are considerably inaccurate in forward prediction.
Interpretable physics discovery. To showcase the interpretability of DisentangO, we visualize its latent variables through latent traversal based on a randomly picked loading field as input in DisentangO-2, as displayed in Figure 6 of Appendix B.3. One can clearly see that the digit changes from “6” to “0” and then “2” from the top left moving down, and from “6” to “1” and then “7” moving to the right. Other digits are visible as well such as “7”, “9”, “4” and “8” in the right-most column. This corresponds well to the distribution of the latent clustering in Figure 3. More discussions on the interpretability of DisentangO-15 can be found in Figures 7 and 8 in Appendix B.3.
4.3 Unsupervised heterogeneous material learning
We now demonstrate the applicability of DisentangO in unsupervised learning in the context of learning synthetic human tissues that exhibit highly organized structures, with collagen fiber arrangements varying spatially. In this subject, it is critical to understand the underlying low-dimensional disentangled properties in the hidden latent space of this complex, high-dimensional microstructure, as inferred from experimental mechanical measurements. We generate two datasets, each containing 500 specimens and 100 loading/displacement pairs. The first dataset features variations in fiber orientation distributions using a Gaussian Random Field (GRF) and the second differs fiber angles in two segmented regions, separated by a centerline with a randomly rotated orientation.
Figure 4: Latent traversal of DisentangO-3 in unsupervised heterogeneous material learning, where the three latent dimensions control the border rotation between the two segments (top), the relative fiber orientation between the two segments (middle), and the fiber orientation of the top segment (bottom), respectively. Legend indicates fiber orientation ranging from 0 to $\pi$.
Table 3: Test errors and the number of trainable parameters in experiment 3. Models are abbreviated (0=DisentangO-2, 1=DisentangO-5, 2=DisentangO-10, 3=DisentangO-15, 4=DisentangO-30, 5=VAE, 6=$\beta$-VAE, 7=MetaNO). Bold number highlights the best method that can handle both forward and inverse settings.
Models #param (M) Error
$\beta_{d}=1$ $\beta_{d}=100$
0 0.63 26.33% 25.18%
1 0.94 17.51% 15.75%
2 1.46 10.76% 10.01%
3 1.98 7.11% 7.02%
4 3.55 5.33% 5.28%
5 3.55 61.10% -
6 3.55 57.04% -
7 0.32 2.67% -
Figure 5: Unsupervised MI score against $\beta_{kl}$ with DisentangO-2 in heterogeneous material learning.
We report in Table 3 our experimental results of DisentangO with different latent dimensions and data loss strength $\beta_{d}$, along with comparisons with the baseline model results. Consistent with the findings in the second experiment, increasing $\beta_{d}$ results in a boost in accuracy by comparing the rows in the table. The model’s predictive performance also converges to the optimal bound of MetaNO as we increase the latent dimension from 2 to 30, where the model’s prediction error improves from 25.18% to 5.28%. This significantly outperforms the best baseline model by 90.7%. On the other hand, the effect of the data loss term on disentanglement is further proved in Figure 5, where increasing $\beta_{d}$ leads to a decrease in MI between the latent factors, thus encouraging disentanglement. Lastly, we interpret the mechanism of the learned latent factors in DisentangO-3 via learning a mapping between the learned latent factors and the underlying material microstructure and subsequently performing latent traversal in each dimension. The results are shown in Figure 4, where the three latent factors manifest control on the border rotation between the two segments, the relative fiber orientation between the two segments, and the fiber orientation of the top segment, respectively.
5 Conclusion
We present DisentangO, a novel hyper-neural operator architecture designed to disentangle latent physical factors embedded within black-box neural operator parameters. DisentangO leverages a hierarchical neural operator architecture that extracts varying parameters of the governing PDE through a task-wise adaptive layer, then further disentangles these variations into distinct latent factors. By learning these disentangled representations, DisentangO not only enhances physical interpretability but also enables more robust generalization across diverse physical systems. We perform empirical evaluations across supervised, semi-supervised, and unsupervised learning contexts.
Limitations: Due to limited computational resource, our experiments are focused on learning a small-to-medium number ($<500$). Expanding the coverage to encompass a broader range of physical systems and enabling learning across more diverse parameters would be beneficial.
Acknowledgments and Disclosure of Funding
L. Zhang would like to acknowledge support by the AFOSR grant FA9550-22-1-0197. Y. Yu would like to acknowledge the support by the National Science Foundation under award DMS-2436624 and the support from the National Institute of Health under award 1R01GM157589-01. Portions of this research were conducted on Lehigh University’s Research Computing infrastructure partially supported by NSF Award 2019035.
Appendix A Proofs of identifiability analysis
We first provide the proof of Theorem 1:
Proof: With equation 3.15, we have:
$p_{\mathcal{G}[\bm{f};\theta]|\bm{f}}=p_{\hat{\mathcal{G}}[\bm{f};\hat{\theta}]|\bm{f}}\Leftrightarrow p_{\mathcal{G}_{f}(\theta)|\bm{f}}=p_{\hat{\mathcal{G}}_{f}(\hat{\theta})|\bm{f}}\Leftrightarrow p_{\theta|\bm{f}}=p_{\mathcal{G}_{f}^{-1}\circ\hat{\mathcal{G}}_{f}(\hat{\theta})|\bm{f}}.$
Note that the parameter $\theta$ varies with the change of $\bm{b}$. Per the data generating process in equation 3.6, the distribution of $\bm{b}$ is invariant to $\bm{f}$. Therefore, the distribution of $\theta$ is also invariant to $\bm{f}$:
$p_{\theta}=p_{\theta|\bm{f}}=p_{\mathcal{G}_{f}^{-1}\circ\hat{\mathcal{G}}_{f}(\hat{\theta})|\bm{f}}=p_{\mathcal{G}_{f}^{-1}\circ\hat{\mathcal{G}}_{f}(\hat{\theta})},\;\forall\bm{f}\in\mathcal{F}.$
Denoting $r:={\mathcal{G}}_{f}\circ\hat{\mathcal{G}}^{-1}_{f}$, it is the transformation between the true $\theta$ and the estimated one, and it is invertible and invariant with respect to $\bm{f}$.
We proceed to derive the relation between $\bm{z}$ and $\hat{\bm{z}}$: since $\theta=r(\hat{\theta})$, with the invertibility assumption $\theta=\mathcal{H}^{-1}(\bm{z})$ and $\hat{\theta}=\hat{\mathcal{H}}^{-1}(\hat{\bm{z}})$, we obtain:
$\bm{z}=\mathcal{H}(\theta)=\mathcal{H}\circ r(\hat{\theta})=\mathcal{H}\circ r\circ\hat{\mathcal{H}}^{-1}(\hat{\bm{z}}).$
Denoting $h:=\mathcal{H}\circ r\circ\hat{\mathcal{H}}^{-1}$, it is the transformation between the true latent variable and the estimated one, and it is invertible because $r$, $\mathcal{H}$ and $\hat{\mathcal{H}}$ are all invertible.∎
We now show the proof of Theorem 2.
Proof: With the independence relation assumption, we have
$p_{\bm{z}|\bm{b}}(\bm{z}|\bm{b})=\prod_{i}p_{z_{i}|\bm{b}}(z_{i}),\;p_{\hat{\bm{z}}|\bm{b}}(\hat{\bm{z}}|\bm{b})=\prod_{i}p_{\hat{z}_{i}|\bm{b}}(\hat{z}_{i})\text{ .}$
Denoting $\hat{q}_{i}:=\log p_{\hat{z}_{i}|\bm{b}}$, it yields:
$\log p_{\bm{z}|\bm{b}}(\bm{z}|\bm{b})=\sum_{i}q_{i}(z_{i},\bm{b}),\;\log p_{\hat{\bm{z}}|\bm{b}}(\hat{\bm{z}}|\bm{b})=\sum_{i}\hat{q}_{i}(\hat{z}_{i},\bm{b})\text{ .}$
With the change of variables we have
$p_{\bm{z}|\bm{b}}=p_{h({\hat{\bm{z}}})|\bm{b}}=p_{{\hat{\bm{z}}}|\bm{b}}\cdot|J_{h^{-1}}|\Leftrightarrow\sum_{i}q_{i}(z_{i},\bm{b})+\log|J_{h}|=\sum_{i}\hat{q}_{i}(\hat{z}_{i},\bm{b})\text{ ,}$
where $|J_{h^{-1}}|$ stands for the absolute value of the Jacobian matrix determinant of $h^{-1}$. Differentiating the above equation twice with respect to $\hat{z}_{k}$ and $\hat{z}_{q}$, $k\neq q$, yields
$\sum_{i}\left(\dfrac{\partial^{2}q_{i}(z_{i},\bm{b})}{\partial z_{i}^{2}}\dfrac{\partial z_{i}}{\partial\hat{z}_{k}}\dfrac{\partial z_{i}}{\partial\hat{z}_{q}}+\dfrac{\partial q_{i}(z_{i},\bm{b})}{\partial z_{i}}\dfrac{\partial^{2}z_{i}}{\partial\hat{z}_{k}\partial\hat{z}_{q}}\right)+\dfrac{\partial^{2}\log|J_{h}|}{\partial\hat{z}_{k}\partial\hat{z}_{q}}=0\text{ .}$ (A.1)
To show the identifiability, one can rewrite the Jacobian $J_{h}$ as:
$J_{h}=\left[\dfrac{\partial\bm{z}}{\partial\hat{\bm{z}}}\right].$
The invertibility results shown in Theorem 1 indicates that it is full rank. Next, we will use the linear independence assumption to show that there exists one and only one non-zero component in each row of $\dfrac{\partial\bm{z}}{\partial\hat{\bm{z}}}$.
Taking $\bm{b}=\bm{b}^{0},\cdots,\bm{b}^{2d_{z}}$ in equation A.1 and subtracting them from each other, we have
$\sum_{i=1}^{d_{z}}\left(\left(\dfrac{\partial^{2}q_{i}(z_{i},\bm{b}^{j})}{\partial z_{i}^{2}}-\dfrac{\partial^{2}q_{i}(z_{i},\bm{b}^{0})}{\partial z_{i}^{2}}\right)\dfrac{\partial z_{i}}{\partial\hat{z}_{k}}\dfrac{\partial z_{i}}{\partial\hat{z}_{q}}+\left(\dfrac{\partial q_{i}(z_{i},\bm{b}^{j})}{\partial z_{i}}-\dfrac{\partial q_{i}(z_{i},\bm{b}^{0})}{\partial z_{i}}\right)\dfrac{\partial^{2}z_{i}}{\partial\hat{z}_{k}\partial\hat{z}_{q}}\right)=0\text{ ,}$
where $j=1,\cdots,2d_{z}$. With the linear independence condition for $w$, this is a $2d_{z}\times 2d_{z}$ linear system, and therefore the only solution is
$\dfrac{\partial z_{i}}{\partial\hat{z}_{k}}\dfrac{\partial z_{i}}{\partial\hat{z}_{q}}=0\text{ ,}\quad\;\dfrac{\partial^{2}z_{i}}{\partial\hat{z}_{k}\partial\hat{z}_{q}}=0\text{ ,}$
for $i=1,\cdots,d_{z}$. The first part implies that, for the $i-$th row of the Jacobian matrix $J_{h}$, we have $\dfrac{\partial z_{i}}{\partial\hat{z}_{k}}\neq 0$ for at most one element $k\in\{1,\cdots,d_{z}\}$, hence $\bm{z}$ is identifiable up to permutation and component-wise invertible transformation. ∎
Appendix B Additional experimental details, results and discussion
Herein, we provide additional details in training and baseline models, as well as more results as a supplement of Section 4.
B.1 Training details
In all experiments, we adopt the Adam optimizer for optimization and use a 16-layer DisentangO model. For fair comparison across different models, we tune the hyperparameters, including the learning rates, the decay rates, and the regularization parameters, to minimize the validation loss. Experiments are conducted on a single NVIDIA Tesla A100 GPU with 40 GB memory. A pseudo algorithm of all three scenarios is summarized in Algorithm 1.
Algorithm 1 A pseudo algorithm of DisentangO.
1:Denote data reconstruction loss $L_{data}$, task-wise NO parameter reconstruction loss $L_{recon}$, KL loss $L_{KL}$ and semi-supervision loss $L_{semi}$ as:
$L_{data}=$ $\dfrac{1}{S}\sum_{\eta=1}^{S}\left[-E(\log(p(\bm{u}|\bm{f},\theta^{\eta})))\right]\text{ ,}\quad L_{recon}=\dfrac{1}{S}\sum_{\eta=1}^{S}\left[-E_{q(\bm{z}^{\eta}|\theta^{\eta})}\log p(\theta^{\eta}|\bm{z}^{\eta})\right]\text{ ,}$
$L_{KL}=$ $\dfrac{1}{S}\sum_{\eta=1}^{S}\left[D_{KL}\left(q(\bm{z}^{\eta}|\theta^{\eta})||p(\bm{z}^{\eta})\right)\right]\text{ ,}\quad L_{semi}=L_{c}(\bm{z}^{\eta},c(\bm{b}^{\eta}))\text{ .}$
2:SC1: Unsupervised/SC3: Supervised
3:The total loss is comprised of the data loss, the NO parameter reconstruction loss, and the KL loss: $L_{loss}=\beta_{d}L_{data}+L_{recon}+\beta_{KL}L_{KL}.$
4:SC2: Semi-supervised
5:The total loss is comprised of the data loss, the NO parameter reconstruction loss, the KL loss, and a semi-supervised loss: $L_{loss}=\beta_{d}L_{data}+L_{recon}+\beta_{KL}L_{KL}+\beta_{cls}L_{semi}.$
B.2 Supervised forward and inverse PDE learning
The parameter of each baseline is given in the following, where the parameter choice of each model is selected by tuning the number of layers and the width (channel dimension), keeping the total number of parameters on the same magnitude.
  • MetaNO: We use a 16-layer IFNO model with the lifting layer as the adaptive layer. We keep the total number of parameters in MetaNO the same as the number of parameters used in forward PDE learning in DisentangO.
  • NIO: We closely follow the setup in Molinaro et al. (2023) and use four convolution blocks as the encoder for the branch net and a fully connected neural network with two hidden layers of 256 neurons as the trunk net, with the number of basis functions set to 50. For the FNO part, we use one Fourier layer with width 32 and modes 8, as suggested in Molinaro et al. (2023).
  • FNO: We adopt a 4-layer FNO with width 26 and modes 8. For the forward model, in addition to the loading field and coordinates as input, we also concatenate the ground-truth material properties to form the final input. For the inverse model, we employ an iterative gradient-based optimization to solve for the optimal material parameters for physics discovery.
  • VAE: We use a 2-layer MLP of size [1681, 136, 30] as the encoder, and another 2-layer MLP of size [30, 136, 1681] as the decoder, with the size of the bottleneck layer being 30.
  • $\beta$-VAE: The parameter choice of the $\beta$-VAE baseline is the same as the VAE baseline, except that we tune the $\beta$ hyperparameter.
B.3 Mechanical MNIST benchmark
Figure 6: MMNIST latent traversal based on a randomly picked loading field as DisentangO-2 input.
Herein, we demonstrate the latent traversal based on a randomly picked loading field as DisentangO-2 input. One can clearly see that the digit changes from “6” to “0” and then “2” from the top left moving down, and from “6” to “1” and then “7” moving to the right. Other digits are visible as well such as “7”, “9”, “4” and “8” in the right-most column. This corresponds well to the distribution of the latent clustering in Figure 3. We also provide two exemplary MMNIST latent interpretations of the learned DisentangO-15 in Figures 7 and 8.
Figure 7: Exemplary MMNIST latent interpretation of DisentangO-15: dimension 10 encodes the information for digit ‘8’, as is evidenced by the fact that all $y<-0.4$ regions on the scatterplots contain only digit ‘8’.
Figure 8: Exemplary MMNIST latent interpretation of DisentangO-15: dimension 12 encodes the information for digit ‘3’, as is evidenced by the fact that all $y<-0.45$ regions on the scatterplots contain only digit ‘3’.
Appendix C Data Generation
C.1 Example 1 - Supervised forward and inverse PDE learning
we consider the constitutive modeling of anisotropic fiber-reinforced hyperelastic materials governed by the Holzapfel–Gasser–Ogden (HGO) model, whose strain energy density function can be written as:
$\eta$ $=\frac{E}{4(1+\nu)}(\bar{I}_{1}-2)-\frac{E}{2(1+\nu)}\ln J$
$+\frac{k_{1}}{2k_{2}}(e^{k_{2}\langle S(\alpha)\rangle^{2}}+e^{k_{2}\langle S(-\alpha)\rangle^{2}}-2)+\frac{E}{6(1-2\nu)}(\frac{J^{2}-1}{2}-\ln J)\text{ ,}$ (C.1)
[…]
Figure 9: Illustration of the HGO data, loading and displacement pairs of three materials $\bm{b}^{\eta}$, $\eta=1,2,3$. Top: three instances of different loadings $T_{y}(x)$, which corresponds to different $\bm{f}_{i}$. Bottom: corresponding displacement solutions $\bm{u}^{\eta}_{i}$, illustrating the impacts of system ($\bm{b}^{\eta}$) variability in solution operators.
C.2 Example 2 - Mechanical MNIST benchmark
Mechanical MNIST is a benchmark dataset of heterogeneous material undergoing large deformation, modeled by the Neo-Hookean material with a varying modulus converted from the MNIST bitmap images (Lejeune, 2020). It contains 70,000 heterogeneous material specimens, and each specimen is governed by the Neo-Hookean material with a varying modulus converted from the MNIST bitmap images. We illustrate samples from the MMNIST dataset in Figure 10, including the underlying microstructure, two randomly picked loading fields, and the corresponding displacement fields.
Figure 10: Illustration of exemplar MMNIST samples in the semi-supervised scenario. (a): material parameter field corresponding to different $\bm{b}^{\eta}$. (b): displacement fields (second row) $\bm{u}^{\eta}_{38}$ corresponding to the same loading field (first row) $\bm{f}_{38}$. (c): displacement fields (second row) $\bm{u}^{\eta}_{138}$ corresponding to the same loading field (first row) $\bm{f}_{138}$.
C.3 Example 3 - Unsupervised heterogeneous material learning
We generate two sets of datasets in this case, varying the material microstructure in the following two ways.
Varying fiber orientation distribution. We generate the samples by controlling the predefined parameters of Gaussian Random Field (GRF). With the GRF sharpened by the thresholding, the values to binary field represent two distinct fiber orientations. The binary field is smoothed using a windowed convolution. To address boundary conditions, the matrix is padded with replicated edge values, ensuring that the convolution works uniformly across the entire grid. After the smoothing process, the padded sections are removed, and the remaining field is used to construct the fiber field. We use two fixed fiber angles $\frac{\pi}{3},\frac{2\pi}{3}$ for the corresponding binary field. We generate 300 material sets, each with 500 loading/displacement pairs, and divide these into training, validation, and test sets in a 200/50/50 split. Exemplar Samples from this dataset are illustrated in Figure 11. These samples demonstrate the variability of $\bm{b}$, which is of critical for the latent variable identifiability in unsupervised learning settings, as proved in Theorem 2.
Varying fiber orientation magnitude and segmentation line rotation. Instead of controlling the fiber orientation angles on the binary field as two constant values, we sample the orientation distribution consisting of two segments with orientations $\alpha_{1}$ and $\alpha_{2}$ on each side, respectively, separated by a line passing through the center. The values of $\alpha_{1}$ and $\alpha_{2}$ are independently sampled from a uniform distribution over $[0,2\pi]$, and the centerline’s rotation is sampled from $[0,\pi]$. We generate 300 material sets, each with 500 loading/displacement pairs, and divide these into training, validation, and test sets in a 200/50/50 split.
Figure 11: Illustration of fiber orientation magnitude and segmentation line rotation. Upper: three different loading instances of $\bm{f}_{i}$. Bottom left: three different microstructure instances of $\bm{b}^{\eta}$. Bottom right: corresponding solution fields $\bm{u}^{\eta}_{i}$.
Loading and displacement pairs for one microstructure. After generating the specimens with varying fiber orientations, we feed the $\bm{b}^{\eta}(x)$ as the $\alpha(x)$ to the HGO model, and keep the material property set $\{E,\nu,k_{1},k_{2}\}$ as constant. For each microstructure, we randomly sample loading and displacement pairs for each microstructure from the previous step.
The loading in this example is taken as the body load, $\bm{f}(\bm{x})$. Each instance is generated as the restriction of a 2D random field, $\phi(\bm{x})=\mathcal{F}^{-1}(\gamma^{1/2}\mathcal{F}(\Gamma))(\bm{x})$. Here, $\Gamma(\bm{x})$ is a Gaussian white noise random field on $\mathbb{R}^{2}$, $\gamma=(w_{1}^{2}+w^{2}_{2})^{-\frac{5}{4}}$ represents a correlation function, $w_{1}$ and $w_{2}$ are the wave numbers on $x$ and $y$ directions, respectively, and $\mathcal{F}$ and $\mathcal{F}^{-1}$ denote the Fourier transform and its inverse, respectively. This random field is anticipated to have a zero mean and covariance operator $C=(-\Delta)^{-2.5}$, with $\Delta$ being the Laplacian with periodic boundary condition on $[0,2]^{2}$, and we then further restrict it to $\Omega$. For the detailed implementation of Gaussian random field sample generation, we refer interested readers to Lang and Potthoff (2011a). Then, for each sampled loading $\bm{f}_{i}(\bm{x})$ and microstructure field $\bm{b}^{\eta}(\bm{x})$, we solve for the displacement field $\bm{u}^{\eta}_{i}(\bm{x})$ on the entire domain.

### v2

@@
 Abstract
-Neural operators (NOs) have demonstrated remarkable success in learning mappings between function spaces, serving as efficient approximators for the forward solutions of complex physical systems governed by partial differential equations (PDEs).
-However, while effective as black-box solvers, they offer limited insight into the underlying physical mechanism, due to the lack of interpretable representations of the physical parameters that drive the system.
-To tackle this challenge, we propose a new paradigm for learning disentangled representations from neural operator parameters, thereby effectively solving an inverse problem.
-Specifically, we introduce DisentangO, a novel hyper-neural operator architecture designed to unveil and disentangle the latent physical factors of variation embedded within the black-box neural operator parameters.
-At the core of DisentangO is a multi-task neural operator architecture that distills the varying parameters of the governing PDE through a task-wise adaptive layer, coupled with a hierarchical variational autoencoder that disentangles these variations into identifiable latent factors.
-By learning these disentangled representations, our model not only enhances physical interpretability but also enables more robust generalization across diverse physical systems.
-Empirical evaluations across supervised, semi-supervised, and unsupervised learning contexts show that DisentangO effectively extracts meaningful and interpretable latent features, bridging the divide between predictive performance and physical understanding in neural operator frameworks.
+Neural operators (NOs) excel at learning mappings between function spaces, serving as efficient forward solution approximators for PDE-governed systems.
+However, as black-box solvers, they offer limited insight into the underlying physical mechanism, due to the lack of interpretable representations of the physical parameters that drive the system.
+To tackle this challenge, we propose a new paradigm for learning disentangled representations from NO parameters, thereby effectively solving an inverse problem.
+Specifically, we introduce DisentangO, a novel hyper-neural operator architecture designed to unveil and disentangle latent physical factors of variation embedded within the black-box neural operator parameters.
+At the core of DisentangO is a multi-task NO architecture that distills the varying parameters of the governing PDE through a task-wise adaptive layer, alongside a variational autoencoder that disentangles these variations into identifiable latent factors.
+By learning these disentangled representations, DisentangO not only enhances physical interpretability but also enables more robust generalization across diverse systems.
+Empirical evaluations across supervised, semi-supervised, and unsupervised learning contexts show that DisentangO effectively extracts meaningful and interpretable latent features, bridging the gap between predictive performance and physical understanding in neural operator frameworks.
+Our code and data accompanying this paper are available at https://github.com/ningliu-iga/DisentangO.
 1 Introduction
 Interpretability in machine learning (ML) refers to the ability to understand and explain how models make decisions (Rudin et al., 2022; Molnar, 2020).
-As ML systems grow more complex, especially with the use of deep learning and ensemble methods (Sagi and Rokach, 2018), the reasoning behind their predictions can become opaque.
+As ML systems grow complex, especially with the use of deep learning and ensemble methods (Sagi & Rokach, 2018), the reasoning behind their predictions can become opaque.
 Interpretability addresses this challenge by making the models more transparent, enabling users to trust the outcomes, detect biases, and identify potential flaws.
@@
 In physics, where models endeavor to capture the governing laws and physical principles, understanding how a model arrives at its predictions is vital for verifying that it aligns with known scientific theories.
-Interpretable models also allow researchers to gain insights into the learned relationships, ensuring that the predictions are physically meaningful and consistent with real-world behaviors.
-This transparency is key for advancing scientific discovery, validating results, and enhancing trust in models used for complex physical systems.
+This transparency is key for advancing scientific discovery, validating results, and enhancing trust in models for complex physical systems.
 Discovering interpretable representations of physical parameters in learning physical systems is challenging due to the intricate nature of real-world phenomena and the often implicit relationships between variables.
-In physics, parameters like force, energy, and velocity are governed by well-established laws, and extracting these parameters in a way that aligns with physical intuition requires models that go beyond mere pattern recognition.
-Traditional ML models may fit the data but fail to represent these parameters in a physically interpretable way.
-To address this, recent developments include the integration of physical constraints into the learning process (Raissi et al., 2019), such as embedding conservation laws (Liu et al., 2023, 2024a), symmetries (Mattheakis et al., 2019), or invariances (Cohen and Welling, 2016) directly into model architectures.
+In physics, quantities like force, energy, and velocity are governed by well-established laws, and extracting them in a way that aligns with physical intuition requires models that go beyond mere pattern recognition.
+Traditional ML models may fit the data but fail to provide a physically interpretable way.
+To address this, recent developments include integrating physical constraints into the learning process (Raissi et al., 2019), such as embedding conservation laws (Liu et al., 2023; 2024a), symmetries (Mattheakis et al., 2019), or invariances (Cohen & Welling, 2016) directly into model architectures.
 Additionally, methods like symbolic regression (Biggio et al., 2021) and sparse modeling (Carroll et al., 2009) aim to discover simple, interpretable expressions that capture the underlying dynamics.
-However, balancing model complexity with interpretability remains a significant challenge, as overly complex models may obscure the true physical relationships, while oversimplified models risk losing critical details of the system’s behavior.
-In this work, we introduce DisentangO, a novel variational hierarchical neural operator architecture to disentangle physical factors of variation from black-box neural operator parameters for solving parametric PDEs.
-Neural operators (NOs) (Li et al., 2020a, c) learn mappings between infinite-dimensional function spaces in the form of integral operators, making them powerful tools for discovering continuum physical laws by manifesting the mappings between spatial and/or spatiotemporal data; see You et al.
+However, balancing model expressivity with interpretability remains a significant challenge, as overly complex models may obscure the true physical relationships, while oversimplified models risk losing critical details of the system’s behavior.
+We introduce DisentangO, a novel variational hyper-neural operator architecture to disentangle physical factors of variation from black-box neural operator parameters for solving parametric PDEs.
+Neural operators (NOs) (Li et al., 2020a; c) learn mappings between infinite-dimensional function spaces in the form of integral operators, making them powerful tools for discovering continuum physical laws by manifesting the mappings between spatial and/or spatiotemporal data; see You et al.
 (2022a); Liu et al.
-(2024a, b, 2023); Ong et al.
+(2024a; b; 2023); Ong et al.
 (2022); Cao (2021); Lu et al.
-(2019, 2021); Goswami et al.
+(2019; 2021); Goswami et al.
 (2022); Gupta et al.
 (2021) and references therein.
-However, most NOs play the role of efficient forward surrogates for the underlying physical system under a supervised learning setting.
-As a result, they act as black-box universal approximators for a single physical system, and are lack of interpretability with respect to the underlying physical laws.
-In contrast, the key innovation of DisentangO lies in its use of a hierarchical architecture that distills the varying physical parameters of the governing PDE from multiple physical systems through an unsupervised learning setting.
-The distilled variations are further disentangled into distinct physical factors to enhance physical understanding and promote robust generalization.
+However, most NOs serve as efficient forward surrogates for the underlying physical system under a supervised learning setting.
+As a result, they act as black-box universal approximators for a single physical system governed by a fixed set of PDE parameters, and lack interpretability with respect to the underlying physical laws.
+In contrast, the key innovation of DisentangO lies in the use of a hypernetwork architecture that distills the varying physical parameters of the governing PDE from multiple physical systems through an unsupervised learning setting.
+The distilled variables are further disentangled into distinct physical factors to enhance physical understanding and promote robust generalization.
 Consequently, DisentangO effectively extracts meaningful and interpretable physical features, thereby simultaneously solving both the forward and inverse problems.
@@
   • We bridge the divide between predictive accuracy and physical interpretability, and introduce a new paradigm that simultaneously performs physics modeling (i.e., as a forward PDE solver) and governing physical mechanism discovery (i.e., as an inverse PDE solver).
-  • We propose a novel variational hierarchical neural operator architecture, which we coin DisentangO.
+  • We propose a novel variational hyper-neural operator architecture, which we coin DisentangO.
 DisentangO extracts the key physical factors of variation from black-box neural operator parameters of multiple physical systems.
-These factors are then disentangled into distinct latent factors that enhance physical interpretation and robust generalization.
+These factors are then disentangled into distinct latent factors that enhance physical interpretation and generalization across different physical systems.
   • We provide theoretical analysis on the component-wise identifiability of the true generative factors in physics modeling: by learning from multiple physical systems, the variability of hidden physical states in these systems promotes identifiability.
-  • We explore the practical utility of disentanglement rather than pursuing it as an end goal, and perform experiments across a broad range of settings including supervised, semi-supervised, and unsupervised learning.
-Our results show that DisentangO effectively extracts meaningful and interpretable physical features for enhanced interpretability.
+  • We explore the practical utility of disentanglement and perform experiments across a broad range of settings including supervised, semi-supervised, and unsupervised learning.
+Results show that DisentangO effectively extracts meaningful physical features.
 2 Background and related work
 Neural operators.
-Learning complex physical systems directly from data is crucial in many scientific and engineering applications (Ghaboussi et al., 1998; Liu et al., 2024c; Ghaboussi et al., 1991; Carleo et al., 2019; Karniadakis et al., 2021; Zhang et al., 2018; Cai et al., 2022; Pfau et al., 2020; He et al., 2021; Jafarzadeh et al., 2024).
-Often, the governing laws of these systems are unknown and must be inferred from data through learning models that are both resolution-invariant to ensure consistent performance across various levels of discretizations and interpretable to domain experts.
-Neural operators (NOs) address the first aim by learning mappings between infinite-dimensional function spaces (Li et al., 2020a, b, c; You et al., 2022a; Ong et al., 2022; Cao, 2021; Lu et al., 2019, 2021; Goswami et al., 2022; Gupta et al., 2021).
-By capturing the relationship between spatial and spatio-temporal data, NOs offer a promising approach to discovering continuum physical surrogates, enabling more accurate and consistent predictions of the underlying responses of complex systems.
-Nevertheless, NOs lack an interpretable representation of the physical states driving the system, and as a result offer limited insight into the underlying physical mechanisms.
+Learning complex physical systems from data is essential in many scientific applications (Carleo et al., 2019; Liu et al., 2024c; Karniadakis et al., 2021; Zhang et al., 2018; Cai et al., 2022; Pfau et al., 2020; He et al., 2021; Jafarzadeh et al., 2024).
+When governing laws are unknown, models must be both resolution-invariant for consistent performance across discretizations and interpretable for domain experts.
+Neural operators (NOs) achieve the former by learning mappings between infinite-dimensional function spaces (Li et al., 2020a; b; c; You et al., 2022a; Ong et al., 2022; Cao, 2021; Lu et al., 2019; 2021; Goswami et al., 2022; Gupta et al., 2021), enabling accurate and consistent predictions of continuum physical surrogates.
+However, NOs lack interpretable representations of physical states, limiting their ability to reveal underlying physical mechanisms.
+Hypernetworks.
+Hypernetworks (Ha et al., 2016; Chauhan et al., 2024) are a class of neural network (NN) architectures that use one NN to generate weights for another NN, both trained in an end-to-end manner.
+This allows soft weight sharing across tasks, benefiting transfer learning and dynamic information sharing (Chauhan et al., 2023).
+Hypernetworks can also be employed as a versatile technique in existing NN architectures.
+For instance, in Nguyen et al.
+; Oh & Peng (2022), a hypernetwork is employed to generate parameters for a VAE model and enable multi-task learning.
+Similarly, Lee et al.
+(2023) integrates the hypernetwork architecture with neural operators.
+However, none of the existing work discusses the capability of hypernetworks in hidden physics discovery.
 Forward and inverse PDE learning.
-Existing NOs focus on providing efficient surrogates for the underlying physical systems as forward PDE solvers.
-They are often employed as black-box universal approximators but lack interpretability of the underlying physical principles.
-On the other hand, several deep learning methods have been developed as inverse PDE solvers (Fan and Ying, 2023; Molinaro et al., 2023; Jiang et al., 2022; Chen et al., 2023), aiming to reconstruct the PDE parameters from solution data.
-However, the inverse problem is typically more challenging due to its ill-posedness.
-To tackle this pathology, many NOs incorporate prior information via governing PDEs (Yang et al., 2021; Li et al., 2021), regularizers (Dittmer et al., 2020; Obmann et al., 2020; Ding et al., 2022; Chen et al., 2023), or additional operator structures (Uhlmann, 2009; Lai et al., 2019; Yilmaz, 2001).
-However, all these work rely on prior information on the operator form, which is not realistic in many real-world problems where neither the parameter field nor the model form are available.
-Worth noting is a recent work that leverages the attention mechanism to build a nonlinear kernel map in NO (Yu et al., 2024), with which a data-dependent kernel can be constructed to characterize the inverse mapping from data to the hidden governing mechanism.
-While effective, this approach does not disentangle the learned kernel or extract physically interpretable parameters.
-To the best of our knowledge, our proposed DisentangO is the first that tackles both forward (i.e., physics prediction) and inverse (i.e., physics discovery) PDE learning problems, while simultaneously inferring the governing distinct physical parameters from the learned NO.
+Existing NOs serve as efficient surrogates for forward PDE solving but often act as black-box approximators, lacking interpretability.
+In contrast, deep learning methods for inverse PDE solving (Fan & Ying, 2023; Molinaro et al., 2023; Jiang et al., 2022; Chen et al., 2023) aim to reconstruct PDE parameters from solution data but face challenges due to ill-posedness.
+To address this, many NOs incorporate prior information via governing PDEs (Yang et al., 2021; Li et al., 2021), regularizers (Dittmer et al., 2020; Obmann et al., 2020; Ding et al., 2022; Chen et al., 2023), or structured operators (Lai et al., 2019; Yilmaz, 2001).
+However, these methods assume prior knowledge of the model form, which is often unrealistic.
+A recent approach (Yu et al., 2024) employs attention to construct a data-dependent kernel for inverse mapping but does not disentangle the learned kernel or extract interpretable parameters.
+To our knowledge, DisentangO is the first to tackle both forward (physics prediction) and inverse (physics discovery) PDE learning while simultaneously identifying distinct physical parameters from the learned NO.
 Disentangled representation learning.
-Disentangled representation learning seeks to separate data into distinct, interpretable factors, where each factor captures an independent aspect of variation in the input.
-This area has critical implications for transfer learning, generative modeling, and fairness in AI, as it allows models to generalize across tasks by leveraging isolated features, making them more interpretable and modular.
-Disentangled representation learning has seen significant advances, primarily driven by models like $\beta$-VAE (Higgins et al., 2017), FactorVAE (Kim and Mnih, 2018), and InfoGAN (Chen et al., 2016), who encourage disentanglement by controlling the trade-off between reconstruction accuracy and latent regularization, or introduce mutual information constraints to encourage the discovery of interpretable latent factors.
-Early work in disentanglement largely focus on supervised or semi-supervised approaches (Ridgeway and Mozer, 2018; Shu et al., 2019; Mathieu et al., 2019), requiring labeled data for factors of variation.
-However, more recent advancements have concentrated on unsupervised learning methods (Duan et al., 2019), where disentanglement is achieved without explicit labels.
-Despite these efforts, fully unsupervised disentanglement remains challenging (Locatello et al., 2019), with ongoing research focusing on improving metrics, robustness, and applicability to more complex and high-dimensional data.
-In this context, although a large body of research has been done in disentanglement, its applications remain largely confined to fields like computer vision and robotics, where the latent dimensions typically correspond to human-interpretable (often visually recognizable) features, with very limited exploration in areas such as physical system learning.
-Moreover, existing work focuses on disentangling from data, whereas our work is the first attempt that explores disentanglement from black-box neural network parameters.
-3 DisentangO: disentangled representation learning from neural operator parameters
+Disentangled representation learning separates data into distinct, interpretable factors, each capturing an independent variation.
+It has critical implications for transfer learning, generative modeling, and AI fairness, enabling models to generalize by leveraging isolated features.
+Key advances have been driven by models like $\beta$-VAE (Higgins et al., 2017), FactorVAE (Kim & Mnih, 2018), and InfoGAN (Chen et al., 2016), which promote disentanglement through latent regularization and mutual information constraints.
+While early work focused on supervised or semi-supervised methods (Ridgeway & Mozer, 2018; Shu et al., 2019; Mathieu et al., 2019), recent efforts target unsupervised learning (Duan et al., 2019), though challenges remain (Locatello et al., 2019), with ongoing efforts to improve metrics, robustness, and applicability to complex data.
+Most research has centered on computer vision and robotics, where latent factors have human-interpretable visual meanings, while its exploration in physical system learning remains limited (Lingsch et al., 2024; Tong et al., 2024; Fotiadis et al., 2023).
+Moreover, existing work disentangles from data, whereas our work is the first to disentangle from black-box NN parameters.
+3 DisentangO
+Figure 1: Overview of the DisentangO architecture.
+Each task correspond to a different (hidden) PDE parameter $\bm{b}$.
+For illustration, the same input function $f^{\eta}_{i}$ is shown for multiple tasks to highlight that different parameter fields $b^{\eta}$ can produce different output functions $u_{i}^{\eta}$ under identical input $f$; in practice, $f$ may vary across tasks.
+The task-specific lifting parameters $\theta_{P}^{\eta}$ are encoded and reconstructed through a VAE, and the reconstructed parameters $\hat{\theta}_{P}^{\eta}$ are fed into the iterative Fourier layers to form the task-specific neural operator $G^{\eta}$.
+Loss components are overlaid to indicate where each term in the objective $L_{loss}$ is computed.
 We consider a series of complex systems with different hidden physical parameters:
-$\mathcal{K}_{\bm{b}}[\bm{u}](\bm{x})=\bm{f}(\bm{x})\text{ ,}\quad\bm{x}\in{\Omega}\text{ .}$ (3.1)
-Here, ${\Omega}\subset\mathbb{R}^{s}$ is the domain of interest, $\bm{f}(\bm{x})$ is a function representing the loading on ${\Omega}$, and $\bm{u}(\bm{x})$ is the corresponding solution of this system.
-$\mathcal{K}_{\bm{b}}$ represents the unknown governing law, e.g., balance laws, which is determined by the (possibly unknown and high-dimensional) parameter field $\bm{b}$.
+[…]
+Here, ${\Omega}\subset\mathbb{R}^{s}$ is the domain of interest, $\bm{f}(\bm{x})$ is a function representing the loading on $\Omega$}, $\bm{u}(\bm{x})$ is the corresponding solution of this system.
+$\mathcal{K}_{\bm{b}}$ represents the unknown governing law, e.g., balance laws, determined by the (possibly unknown and high-dimensional) parameter field $\bm{b}$.
 For instance, in a material modeling problem, $\mathcal{K}_{\bm{b}}$ often stands for the constitutive law and $\bm{b}$ can be a vector ($\bm{b}\in\mathbb{R}^{d_{b}}$) representing the homogenized material parameter field or a vector-valued function ($\bm{b}\in L^{\infty}({\Omega};\mathbb{R}^{d_{b}})$) representing the heterogeneous material properties.
@@
 In a forward problem setting, one aims to find the PDE solution when given PDE information, including coefficient functions, boundary conditions, initial conditions, and loading sources.
-That means, given the governing operators $\mathcal{K}$, the parameter (field) $\bm{b}$, and loading field $\bm{f}(\bm{x})$ in equation 3.1, the goal is to solve for the corresponding solution field $\bm{u}(\bm{x})$, through classical PDE solvers (Brenner and Scott, 2007) or data-driven approaches (Lu et al., 2019; Li et al., 2020c; You et al., 2022b).
+That means, given the governing operators $\mathcal{K}$, the parameter (field) $\bm{b}$, and loading field $\bm{f}(\bm{x})$ in equation 3.1, the goal is to solve for the corresponding solution field $\bm{u}(\bm{x})$, through classical PDE solvers (Brenner & Scott, 2007) or data-driven approaches (Lu et al., 2019; Li et al., 2020c).
 As a result, a forward map is constructed:
@@
 In general, a small number of function pairs $(\bm{u},\bm{f})$ from a single system does not suffice in inferring the underlying parameter field $\bm{b}$, making the inverse problem generally non-identifiable (Molinaro et al., 2023).
-Herein, we propose a novel neural architecture to alleviate the curse of ill-posedness without requiring prior knowledge.
-The key ingredients are: 1) the construction of a multi-task NO architecture, which solves both the forward (physics prediction) and inverse (physics discovery) problems simultaneously from multiple systems, and 2) a generative model to disentangle the features from NO parameters that contain critical information of $\bm{b}$.
-3.1 Variational hierarchical neural operator as a multi-task solver
-Figure 1: Illustration of DisentangO’s architecture.
-Formally, we consider $S$ training datasets from different systems, each of which contains $n_{train}$ function pairs $(\bm{u},\bm{f})\in\mathcal{U}\times\mathcal{F}$:
-$\mathcal{D}_{\rm tr}=\{\{(\bm{u}^{\eta}_{i}(\bm{x}),\bm{f}^{\eta}_{i}(\bm{x}))\}_{i=1}^{n_{train}}\}_{\eta=1}^{S}\text{ .}$ (3.4)
-The index $\eta$ corresponds to different hidden parameter field/set $\bm{b}^{\eta}\in\mathcal{B}$.
-Here, $\mathcal{U}\subset L^{2}({\Omega};\mathbb{R}^{d_{u}})$, $\mathcal{F}\subset L^{2}({\Omega};\mathbb{R}^{d_{f}})$ represent the (infinite-dimensional) Banach spaces of the solution function $\bm{u}$ and loading function $\bm{f}$, respectively, and $\mathcal{B}$ is a finite-dimensional manifold where the hidden parameter $\bm{b}$ takes values on.
-We further assume that the data satisfies the following generating process:
-$\bm{b}|\bm{z}=m(\bm{z})\text{ ,}\;\bm{z}\sim\mathbb{P}_{z}\text{ ,}$ (3.5)
-where $\bm{z}\in\mathcal{Z}\subset\mathbb{R}^{d_{z}}$ is the embedding of $\bm{b}$ in the latent space and $m:\mathcal{Z}\rightarrow\mathcal{B}$ is a one-to-one function.
-Then, we assume for each $\bm{b}^{\eta}$ that there exists an underlying parametric PDE system of equation 3.1.
-Hence, each function pair $(\bm{u}_{i}^{\eta},\bm{f}^{\eta}_{i})$ satisfies:
-$\bm{u}|\bm{f}=\mathcal{K}_{\bm{b}}^{-1}[\bm{f}]\text{ ,}\;\bm{f}\sim\mathbb{P}_{f}\text{ ,}$ (3.6)
-where $\mathcal{K}_{\bm{b}}^{-1}$ is the ground-truth solution operator of the parametric PDE system of equation 3.1 with a given parameter $\bm{b}$.
-Our first objective is to construct a surrogate operator of $\mathcal{K}_{\bm{b}}^{-1}:\mathcal{F}\rightarrow\mathcal{U}$, as the forward PDE solver for each system:
-$\mathcal{G}[\bm{f}^{\eta}_{i};\theta^{\eta}](\bm{x})\approx\mathcal{K}_{\bm{b}^{\eta}}^{-1}[\bm{f}_{i}^{\eta}](\bm{x})=\bm{u}_{i}^{\eta}(\bm{x})\text{ ,}$ (3.7)
-where $\theta^{\eta}$ denotes the corresponding trainable parameter set. Then, as the second objective, we construct an inverse map by discovering the hidden parameter field/set from $\theta^{\eta}$:
-$\mathcal{H}(\theta^{\eta};\Theta_{H})\approx\tilde{\bm{z}}^{\eta}\text{ .}$ (3.8)
-Here, $\Theta_{H}$ is the trainable parameters of $\mathcal{H}$, and $\tilde{\bm{z}}^{\eta}\in\mathbb{R}^{d_{z}}$ denotes the learned latent variables, which can be transformed to the ground-truth latent variable $\bm{z}^{\eta}$ via an invertible function $h$.
-The goal of disentanglement is not to discover $\bm{b}^{\eta}$, since some features may not be identifiable.
+Herein, we propose a novel neural architecture to alleviate the curse of ill-posedness without access to the exact governing partial differential equation (PDE).
+The key ingredients are: 1) the construction of a multi-task NO architecture, which solves both the forward (physics prediction) problems simultaneously from multiple PDE systems with different hidden parameters; 2) a generative model to disentangle the key features from NO parameters that contain critical information of $\bm{b}$, as the inverse (physics discovery) problem solver.
+3.1 Variational hyper-neural operator as a multi-task solver
+Notation and data model assumptions.
+We denote by $\bm{f}(\bm{x})$ and $\bm{u}(\bm{x})$ the input and output functions of the NO, respectively.
+Let $p(\cdot)$ denote a probability density function, $\mathbb{E}(\cdot)$ the expectation, and $\left|\left|\cdot\right|\right|$} the $l^{2}$-norm.
+We assume (noisy) observations of both $\bm{u}$ and $\bm{f}$ are available on a common physical domain $\Omega$, where $\bm{u}\in\mathcal{U}\subset L^{2}(\Omega;\mathbb{R}^{d_{u}})$ and $\bm{f}\in\mathcal{F}\subset L^{2}(\Omega;\mathbb{R}^{d_{f}})$.
+Here, $\mathcal{U}$ and $\mathcal{F}$ are the Banach spaces of solution fields and loading fields, respectively.
+Formally, we consider $S$ training datasets $\mathcal{D}^{\eta}$, $\eta=1,\dots,S$, each corresponding to the same PDE equation 3.1 but with a different (hidden) physical parameter $\bm{b}^{\eta}\in\mathcal{B}$, with $\mathcal{B}$ the parameter space.
+In our multi-task learning model, each dataset corresponds to one task.
+We assume the hidden parameter is generated according to:
+$\bm{b}\sim\mathbb{P}_{b}\text{ ,}\;\bm{z}\sim p(\bm{z}|\bm{b})\text{ ,}$ (3.4)
+where $\bm{z}\in\mathcal{Z}\subset\mathbb{R}^{d_{z}}$ is the latent embedding of $\bm{b}$.
+Each dataset contains measurements from $n^{\eta}_{train}$ function pairs $\{(\bm{u}_{i}^{\eta}(\bm{x}),\bm{f}_{i}^{\eta}(\bm{x}))\}_{i=1}^{n^{\eta}_{train}}$.
+Although our method accommodates datasets with varying numbers of function pairs, we use the same number of function pairs in this work for simplicity and denote $n^{\eta}_{train}=n_{train}$.
+While actual observations are gathered on a discrete sensor set $\chi=\{\bm{x}_{j}\}_{j=1}^{\#\chi}$ and inevitably contain observational noise of the solution, we adopt the standard assumption that this noise is additive and i.i.d.
+Letting $\mathcal{G}^{\dagger}[\bm{f};\bm{b}]$ denote the true forward operator of equation 3.1, we write the data model associated with multi-task learning as:
+[…]
+with
+$\bm{f}\sim\mathbb{P}_{f},\,\bm{f}^{\eta}_{i,j}=\bm{f}^{\eta}_{i}(\bm{x}_{j}),\,\bm{u}^{\eta}_{i,j}=\bm{u}^{\eta}_{i}(\bm{x}_{j})+\epsilon_{\eta,i,j}=\mathcal{G}^{\dagger}[\bm{f}^{\eta}_{i};\bm{b}^{\eta}](\bm{x}_{j})+\epsilon_{\eta,i,j},\;\epsilon_{\eta,i,j}\sim\mathcal{N}(0,\varpi^{2}).$ (3.6)
+The forward modeling objective is to learn a surrogate solution operator $\mathcal{G}[\cdot;\theta^{\eta}]$ for $\mathcal{G}^{\dagger}$, where $\theta^{\eta}$ denotes the task-specific NO parameters for the forward surrogate operator in the $\eta$-th task, associated with $\bm{b}^{\eta}$.
+Note that all tasks share the same NO architecture, $\mathcal{G}$, and the surrogate operator for each task depends on the physical system $\eta$ through the task-specific parameter $\theta^{\eta}$.
+For inverse modeling, let $\mathcal{H}^{\dagger}$ denote the (unknown) inverse operator that maps function pairs $(\bm{u},\bm{f})$ to the underlying PDE parameter $\bm{b}$.
+In practice, identifying $\mathcal{G}^{\dagger}$is often impossible, as the available data of $(\bm{u},\bm{f})$ may not contain sufficient information to identify all features of $\bm{b}$.
 For instance, in a Dirichlet boundary condition problem, $\bm{u}^{\eta}_{i}(\bm{x})=0$ for all $\bm{x}\in\partial\Omega$, making $\bm{b}^{\eta}$ not learnable on $\partial\Omega$.
-Hence, a more realistic goal is to discover the underlying mechanism of $\bm{b}$ in the space of identifiability, i.e., $\bm{z}$, or its equivalent transformation $\tilde{\bm{z}}$.
-Note that in the first objective of equation 3.7, one is interested in finding the solution operator $\mathcal{G}^{\eta}$ under a supervised setting, in the form of function-to-function mappings, for all (hidden) parameters $\bm{b}^{\eta}$ in the range of interest; while in the second objective of equation 3.8, the goal is to construct a vector-to-vector mapping from $\theta^{\eta}$ to $\tilde{\bm{z}}^{\eta}$, which is unsupervised due to the hidden physics nature.
-We propose to employ variational autoencoder (VAE) (Kingma and Welling, 2013) as the representation learning approach for the second objective, and a meta-learned NO architecture as a universal solution operator for the first objective.
-The key is to pair these architectures as a hierarchical neural operator, which simultaneously solves both forward and inverse modeling problems.
-We now define the learning architecture. The overall objective is to maximize the log data likelihood:
-$\max E(\log(p(\theta,\mathcal{D})))=\max\left[E(\log(p(\bm{u}|\bm{f},\theta)+\log(p(\theta|\bm{f}))+\log(p(\bm{f}))\right].$
+A more realistic goal is therefore to recover the underlying mechanism of $\bm{b}$ in the space of identifiability, i.e., $\bm{z}$.
+Thus, our second objective is to construct an inverse map by estimating embedding $\tilde{\bm{z}}^{\eta}$ of the hidden parameter $\bm{b}^{\eta}$ from $\theta^{\eta}$:
 […]
+where $\Theta_{H}$ are trainable parameters of $\mathcal{H}$, and $\tilde{\bm{z}}^{\eta}\in\mathbb{R}^{d_{z}}$ denotes the learned latent variables that can be transformed to the ground-truth latent variable $\bm{z}^{\eta}$ via an invertible function $h$.
+The goal of disentanglement is to discover $\tilde{\bm{z}}$, together with the solution operator $\mathcal{G}$.
+The first objective learns the forward operator $\mathcal{G}[\cdot;\theta^{\eta}]$ in a supervised fashion, in the form of function-to-function mappings, for all (hidden) parameters $\bm{b}^{\eta}$ in the range of interest; while the second objective learns the vector-to-vector mapping from $\theta^{\eta}$ to $\tilde{\bm{z}}^{\eta}$, which is unsupervised due to the hidden physics nature.
+We propose to employ a variational autoencoder (VAE) (Kingma & Welling, 2013) as the representation learning approach for the second objective, and a meta-learned NO architecture as a universal solution operator for the first objective***Note that tasks differ in the PDE coefficients (the physical parameters $b^{\eta}$) and correspondingly the NO parameter $\theta^{\eta}$, not in the NO architecture or its input-output structure..
+The key is to pair these architectures as a hyper-neural operator, which simultaneously solves both forward and inverse problems.
+Although the proposed strategy is generic and thus applicable to other multi-task neural operator architectures, to provide a universal architecture of $\mathcal{G}$ for different tasks, we adopt a meta-learning strategy following the meta-learned neural operator (MetaNO) (Zhang et al., 2023).
+MetaNO is developed based on the implicit Fourier neural operator (IFNO), a PDE solution operator with a relatively small number of trainable parameters (You et al., 2022b).
+In MetaNO, task-wise adaptation is applied only to the trainable parameters in the first layer of the NO, whereas all other parameters are shared across tasks.
+For an $L$-layer MetaNO, we write:
 […]
+where $\mathcal{P}$, $\mathcal{Q}$ are shallow-layer MLPs that map a low-dimensional vector to a high-dimensional vector and vice versa, parameterized by $\theta^{\eta}_{P}$ and $\theta_{Q}$, respectively.
+Each intermediate layer, $\mathcal{J}$, is constructed as a mimetic of a fixed-point iteration step and parameterized by $\theta_{J}$.
+Supported by the universal approximator analysis (Zhang et al., 2023), different PDEs share common iterative ($\theta_{J}$) and projection ($\theta_{Q}$) parameters, with all the information about parameter $\bm{b}$ encoded in the task-wise lifting parameters $\theta^{\eta}_{P}$.
+Hence, to provide an inverse map from the task-wise NO parameter $\theta^{\eta}$ to the key features of the PDE parameters, $\tilde{\bm{z}}^{\eta}$, one can construct $\mathcal{H}$ as the mapping from $\theta^{\eta}_{P}$ to $\tilde{\bm{z}}^{\eta}$ using a neural network architecture.
+In this work, we employ a standard MLP:
+[…]
+since all other NO parameters are invariant to the change of $\bm{b}$.
+This construction substantially reduces the degrees of freedom in $\theta^{\eta}$, making the invertibility assumption in the next section feasible.
+To simplify notation, we use $\theta^{\eta}$ to denote $\theta^{\eta}_{P}$ in the subsequent discussion.
+An overview of the forward and inverse operator architecture is provided in Figure 1.
+We now define the learning objective. The overall objective is to maximize the log data likelihood:
+$\max\mathbb{E}(\log p(\theta,\mathcal{D}))=\max\Big[\mathbb{E}\big(\log p(\bm{u}|\bm{f},\theta)+\log p(\theta|\bm{f})+\log p(\bm{f})\big)\Big].$
+Note that $p(\bm{f})$ remains constant over different NO parameter $\theta$, and the assumption in eq.
+3.6 yields $\log(p(\theta|\bm{f}))=\log(p(\theta))$.
+Additionally, the assumption in eq.
+3.4 guarantees that $\theta$ is generated from the latent space over $\bm{z}$, hence $\log(p(\theta))=\log\int_{\bm{z}}p(\theta|\bm{z})p(\bm{z})d\bm{z}$.
+The overall objective then becomes:
+$\max\mathbb{E}(\log(p(\theta,\mathcal{D})))=\max\bigg[\mathbb{E}(\log(p(\bm{u}|\bm{f},\theta)))+\mathbb{E}\Big(\log\int_{\bm{z}}p(\theta|\bm{z})p(\bm{z})d\bm{z}\Big)\bigg].$ (3.10)
 However, the second term in this formulation is generally intractable.
 We use a variational posterior $q(\theta|\bm{z})$ to approximate the actual posterior $p(\theta|\bm{z})$ and maximize the evidence lower bound (ELBO):
-$L_{ELBO}=\dfrac{1}{S}\sum_{\eta=1}^{S}\left[E_{q(\bm{z}^{\eta}|\theta^{\eta})}\log p(\theta^{\eta}|\bm{z}^{\eta})-D_{KL}\left(q(\bm{z}^{\eta}|\theta^{\eta})||p(\bm{z}^{\eta})\right)\right].$
-Putting everything together, we obtain the loss functional:
-$L_{loss}=\dfrac{1}{S}\sum_{\eta=1}^{S}\left[-E(\log(p(\bm{u}|\bm{f},\theta^{\eta})))-E_{q(\bm{z}^{\eta}|\theta^{\eta})}\log p(\theta^{\eta}|\bm{z}^{\eta})+D_{KL}\left(q(\bm{z}^{\eta}|\theta^{\eta})||p(\bm{z}^{\eta})\right)\right].$ (3.10)
-The above formulation naturally lends itself to a hierarchical variational autoencoder (HVAE) architecture (Vahdat and Kautz, 2020).
-The encoder aims to obtain the posterior $q_{\mu_{z},\Sigma_{z}}(\bm{z}^{\eta}|\theta^{\eta})$, which estimates the inverse map $\mathcal{H}$:
-$\hat{\bm{z}}\sim q_{\mu_{z},\Sigma_{z}}(\hat{\bm{z}}^{\eta}|\theta^{\eta})\text{ .}$ (3.11)
-Then, the first decoder, which processes the estimated latent variable $\hat{\bm{z}}$ and reconstructs the corresponding NO parameter $\hat{\theta}$:
-$\hat{\theta}=\hat{g}(\hat{\bm{z}})\text{ .}$ (3.12)
+[…]
+with $D_{KL}(\cdot||\cdot)$ denoting the KL divergence between two distributions. Putting everything together, we obtain the loss functional:
+$L_{loss}=\frac{1}{S}\sum_{\eta=1}^{S}\bigg[-\mathbb{E}(\log(p(\bm{u}|\bm{f},\theta^{\eta})))-\mathbb{E}_{q(\bm{z}^{\eta}|\theta^{\eta})}\log p(\theta^{\eta}|\bm{z}^{\eta})+D_{KL}\left(q(\bm{z}^{\eta}|\theta^{\eta})\|p(\bm{z}^{\eta})\right)\bigg].$ (3.11)
+The above formulation naturally lends itself to a hierarchical variational autoencoder (HVAE) architecture (Vahdat & Kautz, 2020).
+The encoder aims to obtain the posterior $q_{\mu_{z},\Sigma_{z}}(\bm{z}^{\eta}|\theta^{\eta})$ and provide the inverse map $\mathcal{H}$:
+[…]
+Note that although $\mathcal{H}$ represents a deterministic inverse mapping from the NO parameters to the latent variables, in equation 3.12 we approximate this mapping using a probabilistic encoder to account for uncertainty and enable variational inference.
+This follows standard practice in the VAE literature, where deterministic relationships are modeled probabilistically for tractable inference.
+Then, the first decoder $\hat{g}$ processes the estimated latent variable $\hat{\bm{z}}$ and reconstructs the corresponding NO parameter $\hat{\theta}$:
+[…]
 Lastly, the second decoder reconstructs the forward map $\mathcal{G}$, by further taking the estimated $\hat{\theta}$ and the loading function $\bm{f}$ and estimating the output function $\bm{u}$:
-$\hat{\bm{u}}=\hat{\mathcal{G}}[\bm{f};\hat{\theta}]\text{ .}$ (3.13)
-For this purpose, we employ the implicit Fourier neural operator (IFNO)***We also point out that the proposed strategy is generic and hence also applicable to other neural operators., which is a provably universal approximator for PDE solution operators with a relatively small number of trainable parameters (You et al., 2022b).
-To provide a universal $\hat{\mathcal{G}}$ for different domains, a multi-task learning strategy is adopted following Zhang et al.
-(2023), which applies task-wise adaptation only to the first layer of NOs.
-Specifically, an $L$-layer multi-task IFNO has the following form:
-${\mathcal{G}}[\bm{f};{\theta}^{\eta}](\bm{x})={\mathcal{G}}[\bm{f};{\theta}_{P}^{\eta},\theta_{J},\theta_{Q}](\bm{x}):=\mathcal{Q}_{\theta_{Q}}\circ(\mathcal{J}_{\theta_{J}})^{L}\circ\mathcal{P}_{{\theta}^{\eta}_{P}}[\bm{f}](\bm{x})\text{ ,}$ (3.14)
-where $\mathcal{P}$, $\mathcal{Q}$ are shallow-layer MLPs that map a low-dimensional vector into a high-dimensional vector and vice versa, parameterized by $\theta_{Q}$ and ${\theta}^{\eta}_{P}$, respectively.
-Each intermediate layer, $\mathcal{J}$, is constructed as a mimetic of a fixed-point iteration step and parameterized by $\theta_{J}$.
-Supported by the universal approximator analysis in Zhang et al.
-(2023), different PDEs share the common iterative ($\theta_{J}$) and projection ($\theta_{Q}$) parameters, with all the information about parameter $\bm{b}$ embedded in the task-wise lifting parameters $\theta^{\eta}_{P}$.
-Hence, the encoder and first decoder structure only need to encode and reconstruct $\theta^{\eta}_{P}$, since all other NO parameters are invariant to the change of $\bm{z}$.
-This construction substantially reduces the degree of freedoms in $\theta^{\eta}$, making the invertibility assumption in the next section feasible.
-With a slight abuse of notation, in the following contents we use $\theta^{\eta}$ to denote the task-wise parameter, i.e., the lifting layer parameter $\theta^{\eta}_{P}$, and estimate $\bm{z}^{\eta}$ from it.
+[…]
 3.2 Disentangling the underlying mechanism
-The proposed DisentangO aims to identify and disentangle the components of the latent representation $\bm{z}$, which serves as an inverse PDE solver.
-However, to capture the true physical mechanism, a natural question arises: is it really possible to identify the latent variables of interest (i.e., $\bm{z}$) with only observational data $\mathcal{D}_{\rm tr}$?
-We now show that, by learning the model $(p_{\hat{\bm{z}}},\hat{g},\hat{\mathcal{H}},\hat{\mathcal{G}})$ that matches the true marginal data distribution in all domains, we can indeed achieve this identifiability for the generating process proposed in equation 3.5 and equation 3.6.
+[…]
+Assumption 1.
+(Density Smoothness and Positivity) The probability density functions for $\theta$ and $\bm{z}$, which we denote as $p_{\theta}$ and $p_{\bm{z}}$, are both smooth and positive.
+Assumption 2.
+(Invertibility) The task-wise parameter $\theta$ can be generated by $\bm{z}$ through an invertible and smooth function $\mathcal{H}^{-1}$.
+Moreover, for each given $\bm{f}$, we denote $\mathcal{G}_{f}(\theta):=\mathcal{G}[\bm{f};\theta](\bm{x})$ as the operator mapping from $\mathbb{R}^{d_{\theta}}$ to $\mathcal{U}$.
+$\mathcal{G}_{f}$ is also one-to-one with respect to $\theta$.
+Assumption 3.
+(Conditional Independence) Conditioned on $\bm{b}$, each component of $\bm{z}$ is independent of each other: $\log p_{\bm{z}|\bm{b}}(\bm{z}|\bm{b})=\sum_{i=1}^{d_{z}}\log p_{z_{i}|\bm{b}}(z_{i}|\bm{b})$.
+Assumption 4.
+(Linear independence) There exists $2d_{z}+1$ values of $\bm{b}$, such that the $2d_{z}$ vectors $w(\bm{z},\bm{b}^{j})-w(\bm{z},\bm{b}^{0})$ with $j=1,\cdots,2d_{z}$ are linearly independent. Here,
+$w(\bm{z},\bm{b}^{j}):=\Bigg(\dfrac{\partial q_{1}(z_{1},\bm{b}^{j})}{\partial z_{1}},\cdots,\dfrac{\partial q_{d_{z}}(z_{d_{z}},\bm{b}^{j})}{\partial z_{d_{z}}},\dfrac{\partial^{2}q_{1}(z_{1},\bm{b}^{j})}{\partial z_{1}^{2}},\cdots,\dfrac{\partial^{2}q_{d_{z}}(z_{d_{z}},\bm{b}^{j})}{\partial z_{d_{z}}^{2}}\Bigg),\vskip-5.78172pt$ (3.15)
+with $q_{i}(z_{i},\bm{b}):=\log p_{z_{i}|\bm{b}}$.
 First, we show that the latent variable $\bm{z}$ can be identified up to an invertible component-wise transformation: for the true latent variable $\bm{z}\in\mathbb{R}^{d_{z}}$, there exists an invertible function $h:\mathbb{R}^{d_{z}}\rightarrow\mathbb{R}^{d_{z}}$, such that $\hat{\bm{z}}=h(\bm{z})$.
 Theorem 1.
-We follow the data-generating process in equation 3.5 and equation 3.6, and make the following assumptions:
-  • (Density Smoothness and Positivity): $p_{\theta}$ and $p_{\bm{z}}$ are both smooth and positive.
-  • (Invertibility): The task-wise parameter $\theta$ can be generated by $\bm{z}$ through an invertible and smooth function $\mathcal{H}^{-1}$.
-Moreover, for each given $\bm{f}$, we denote $\mathcal{G}_{f}(\theta):=\mathcal{G}[\bm{f};\theta](\bm{x})$ as the operator mapping from $\mathbb{R}^{d_{\theta}}$ to $\mathcal{U}$.
-$\mathcal{G}_{f}$ is also one-to-one with respect to $\theta$.
-Then, by learning $(p_{\hat{\bm{z}}},\hat{g},\hat{\mathcal{H}},\hat{\mathcal{G}})$ to achieve:
+[…]
 $p_{\hat{\bm{u}}|\bm{f}}=p_{\bm{u}|\bm{f}}\text{ ,}$ (3.16)
@@
 Theorem 2.
-We follow the data-generating process in equation 3.5 and equation 3.6, and assume that the Density Smoothness and Positivity condition as well as the Invertibility condition in Theorem 1 are both satisfied, together with the following additional assumptions:
-  • (Conditional Independence): Conditioned on $\bm{b}$, each component of $\bm{z}$ is independent of each other: $\log p_{\bm{z}|\bm{b}}(\bm{z}|\bm{b})=\sum_{i=1}^{d_{z}}\log p_{z_{i}|\bm{b}}(z_{i}|\bm{b})$.
-  • (Linear independence) There exists $12d_{z}+$ 1 values of $\bm{b}$, such that the $2d_{z}$ vectors $w(\bm{z},\bm{b}^{j})-w(\bm{z},\bm{b}^{0})$ with $j=1,\cdots,2d_{z}$ are linearly independent.
-Here,
-    $w(\bm{z},\bm{b}^{j}):=\left(\dfrac{\partial q_{1}(z_{1},\bm{b}^{j})}{\partial z_{1}},\cdots,\dfrac{\partial q_{d_{z}}(z_{d_{z}},\bm{b}^{j})}{\partial z_{d_{z}}},\dfrac{\partial^{2}q_{1}(z_{1},\bm{b}^{j})}{\partial z_{1}^{2}},\cdots,\dfrac{\partial^{2}q_{d_{z}}(z_{d_{z}},\bm{b}^{j})}{\partial z_{d_{z}}^{2}}\right),$ (3.16)
-    with $q_{i}(z_{i},\bm{b}):=\log p_{z_{i}|\bm{b}}$.
-Then, by learning $(p_{\hat{\bm{z}}},\hat{g},\hat{\mathcal{H}},\hat{\mathcal{G}})$ to achieve equation 3.15, $\bm{z}$ is component-wise identifiable.
+[…]
 Proof: Please see Appendix A.
-Intuitively, the linear independence condition in Theorem 2 indicates sufficient variability between physical systems (tasks).
+Discussion on assumptions.
+Intuitively, Assumptions 1-2 are required to guarantee that there exists a smooth and injective mapping between the ground-truth latent embedding $\bm{z}$ and the learned embedding $\hat{\bm{z}}$.
+Assumption 4 indicates sufficient variability across physical systems.
 This is a common assumption in the nonlinear ICA literature for domain adaptation (Hyvarinen et al., 2019; Khemakhem et al., 2020; Kong et al., 2023).
-To our best knowledge, this is the first time the component-wise identifiability results are discussed in the context of multi-task neural operator learning.
-3.3 A generic algorithm for different supervision scenarios
+Further discussions and validation on the assumptions are provided in Appendix A.
+To our best knowledge, it is the first time the component-wise identifiability is discussed in the context of multi-task neural operator learning.
+3.3 A generic algorithm for various supervision cases
+Although DisentangO is primarily designed for the challenging scenario of learning without supervision on $\bm{b}$ nor prior knowledge on the PDE model form in equation 3.1, its methodology is generic and readily can be extended to handle the scenarios with partial or full measurements of $\bm{b}$, which are common in classical inverse PDE benchmark problems.
 In this section, we discuss the practical utility of DisentangO under three scenarios:
-  • (SC1: Unsupervised) No value or label is given for $\bm{b}^{\eta}$ for each task.
+  • (SC1: Supervised) The value of $\bm{b}^{\eta}$ is given.
   • (SC2: Semi-supervised) The value of $\bm{b}^{\eta}$ is not given, but a label $c(\bm{b}^{\eta})$ (e.g., a classification of $\bm{b}^{\eta}$) is given.
-  • (SC3: Supervised) The value of $\bm{b}^{\eta}$ is given.
-The architecture is shown in Figure 1, with a pseudo algorithm provided in Algorithm 1 of Appendix.
-To obtain the posterior $q$ in equation 3.11, we assume each latent variable satisfies a Gaussian distribution of distinct means and diagonal covariance, i.e., $q_{\mu_{z},\Sigma_{z}}(\hat{\bm{z}}|\theta):=\mathcal{N}(\mu_{z}(\theta_{P}),\Sigma^{2}_{z}(\theta_{P}))$, then estimate its mean and covariance using an MLP.
+  • (SC3: Unsupervised) No value or label is given for $\bm{b}^{\eta}$ for each task.
+A pseudo Algorithm 1 is provided in the Appendix.
+To obtain the posterior $q$ in equation 3.12, we assume that each latent variable satisfies a Gaussian distribution of distinct means and diagonal covariance, i.e., $q_{\mu_{z},\Sigma_{z}}(\hat{\bm{z}}|\theta):=\mathcal{N}(\mu_{z}(\theta),\Sigma^{2}_{z}(\theta))$, then estimate its mean and covariance using an MLP.
 As a result, the KL-divergence term in the ELBO admits a closed form:
+$\text{SC2/SC3:}\quad D_{KL}\left(q(\hat{\bm{z}}|\theta)\parallel p(\bm{z})\right)=\frac{1}{2}\sum_{i=1}^{d_{z}}\Big((\Sigma_{z})_{i}^{2}+(\mu_{z})^{2}_{i}-2\log((\Sigma_{z})_{i})-1\Big)\text{.}$ (3.17)
+In the supervised setting, we take $\mu_{z}(\theta_{P}^{\eta})=\bm{b}^{\eta}$, and then the KL divergence term writes:
+$\text{SC1:}\quad D_{KL}\left(q(\hat{\bm{z}}|\theta)\parallel p(\bm{z})\right)=\frac{1}{2}\sum_{i=1}^{d_{z}}\Big((\Sigma_{z})_{i}^{2}+(\mu_{z}-\bm{b})^{2}_{i}-2\log((\Sigma_{z})_{i})-1\Big)\text{.}\vskip-6.50403pt$ (3.18)
+For the first decoder, the likelihood $p(\theta^{\eta}|\bm{z}^{\eta})$ is a factorized Gaussian with mean $\mu_{\theta}$ and covariance $\Sigma_{\theta}$, computed from another MLP.
+By taking as input Monte Carlo samples once for each $\bm{z}$, the reconstruction accuracy term can be approximated as:
 […]
-In the supervised setting, we take $\bm{z}=\bm{b}$, and set $\mu_{z}(\theta_{P}^{\eta})=\bm{b}^{\eta}$, then:
+where $d_{\theta}$} is the dimension of $\theta$, and $c$ is a constant.
+For the second decoder, we parameterize it as a multi-task NO in equation 3.8, and model the discrepancy between $\hat{\mathcal{G}}[\bm{f}^{\eta}_{j};\hat{\theta}^{\eta}](\bm{x}_{k})$ and the ground truth $\bm{u}(\bm{x}_{k})$ as an additive independent unbiased Gaussian random noise $\epsilon$, with
+$\hat{\mathcal{G}}[\bm{f}^{\eta}_{j};\hat{\theta}^{\eta}](\bm{x}_{k})=\bm{u}_{j}^{\eta}(\bm{x}_{k})+\epsilon_{\eta,j,k}\text{, }\epsilon_{\eta,j,k}\sim\mathcal{N}(0,\varpi^{2})\text{,}$
+where $\varpi$ is the standard deviation of the additive noise as defined in the data model.
+In practice, the observational noise is unknown, and we treat $\varpi$ as a tunable hyperparameter.
+Then, with a uniform spatial discretization of size $\Delta x$ in a domain ${\Omega}\subset\mathbb{R}^{d_{u}}$, the log likelihood after eliminating the constant terms can be written as:
 […]
-For the first decoder, the likelihood $p(\theta^{\eta}_{P}|\bm{z}^{\eta})$ is a factorized Gaussian with mean $\mu_{\theta}$ and covariance $\Sigma_{\theta}$, computed from another MLP.
-By taking as input Monte Carlo samples once for each $\bm{z}$, the reconstruction accuracy term can be approximated as:
-$E_{q(\bm{z}|\theta_{P})}\log p(\theta_{P}|\bm{z})\approx-\sum_{i=1}^{d_{\theta}}\left(\dfrac{((\theta_{P})_{i}-(\mu_{\theta})_{i})^{2}}{2(\Sigma_{\theta})^{2}_{i}}+\log(\Sigma_{\theta})_{i}+c\right)\text{ ,}$
-where ${d_{\theta}}$ is the dimension of $\theta_{P}$, and $c$ is a constant.
-For the second decoder, we parameterize it as a multi-task NO in equation 3.14, and model the discrepancy between $\hat{\mathcal{G}}_{[\hat{\theta}^{\eta}_{P},\theta_{I},\theta_{Q}]}[\bm{g}^{\eta}_{j}](\bm{x}_{k})$ and the ground truth $\bm{u}(\bm{x}_{k})$ as an additive independent unbiased Gaussian random noise $\epsilon$, with
-$\hat{\mathcal{G}}_{[\hat{\theta}^{\eta}_{P},\theta_{I},\theta_{Q}]}[\bm{g}^{\eta}_{j}](\bm{x}_{k}))=\bm{u}_{j}^{\eta}(\bm{x}_{k})+\epsilon_{\eta,j,k}\text{ ,}\quad\epsilon_{\eta,j,k}\sim\mathcal{N}(0,\varpi^{2})\text{ ,}$
-where $\varpi$ is a tunable hyperparameter.
-Then, with a uniform spatial discretization of size $\Delta x$ in a domain ${\Omega}\subset\mathbb{R}^{d_{u}}$, the log likelihood after eliminating the constant terms can be written as:
-$\dfrac{1}{2\varpi^{2}\Delta x^{2d_{u}}}\sum_{\eta=1}^{S}\sum_{j=1}^{n_{train}}{\left|\left|\hat{\mathcal{G}}_{[\hat{\theta}^{\eta}_{P},\theta_{I},\theta_{Q}]}[\bm{g}^{\eta}_{j}](\bm{x}_{k})-\bm{u}_{j}^{\eta}(\bm{x}_{k})\right|\right|}_{L^{2}({\Omega})}^{2}.$
-In our empirical tests in Section 4, for the unsupervised and semi-supervised scenarios we choose $q$ as a standard Gaussian distribution following the independence assumption of Theorem 2.
+In our empirical tests in Section 4, for the unsupervised and semi-supervised scenarios, we choose $q$ as a standard Gaussian distribution following the independence assumption of Theorem 2.
 Additionally, to avoid overparameterization, we take $\Sigma_{\theta}=\sigma^{2}_{\theta}\mathbf{I}$, with $\sigma_{\theta}$ being a tunable hyperparameter.
 In this case, the total objective can be further simplified as:
-$L_{loss}=$ $\dfrac{1}{S}\sum_{\eta=1}^{S}\left(\beta_{d}\sum_{j=1}^{n_{train}}{\left|\left|\hat{\mathcal{G}}_{[\hat{\theta}^{\eta}_{P},\theta_{I},\theta_{Q}]}[\bm{g}^{\eta}_{j}]-\bm{u}^{\eta}_{j}\right|\right|}^{2}_{L^{2}({\Omega})}+{\left|\left|\hat{\theta}^{\eta}_{P}-\mu^{\eta}_{\theta}\right|\right|}^{2}+\beta_{KL}{\left|\left|\mu^{\eta}_{\bm{z}}\right|\right|}^{2}\right)\text{ ,}$
+[…]
 for the unsupervised scenario, and
-$L_{loss}=$ $\dfrac{1}{S}\sum_{\eta=1}^{S}\left(\beta_{d}\sum_{j=1}^{n_{train}}{\left|\left|\hat{\mathcal{G}}_{[\hat{\theta}^{\eta}_{P},\theta_{I},\theta_{Q}]}[\bm{g}^{\eta}_{j}]-\bm{u}^{\eta}_{j}\right|\right|}^{2}_{L^{2}({\Omega})}+{\left|\left|\hat{\theta}^{\eta}_{P}-\mu^{\eta}_{\theta}\right|\right|}^{2}+\beta_{KL}{\left|\left|\mu^{\eta}_{\bm{z}}-\bm{b}^{\eta}\right|\right|}^{2}\right)\text{ ,}$
+[…]
 for the fully supervised scenario.
@@
 (2020):
-$L_{loss}=\dfrac{1}{S}\sum_{\eta=1}^{S}$ $\left(\beta_{d}\sum_{j=1}^{n_{train}}{\left|\left|\hat{\mathcal{G}}_{[\hat{\theta}^{\eta}_{P},\theta_{I},\theta_{Q}]}[\bm{g}^{\eta}_{j}]-\bm{u}^{\eta}_{j}\right|\right|}^{2}_{L^{2}({\Omega})}+{\left|\left|\hat{\theta}^{\eta}_{P}-\mu^{\eta}_{\theta}\right|\right|}^{2}\right.+\left.\beta_{KL}{\left|\left|\mu^{\eta}_{\bm{z}}\right|\right|}^{2}+\beta_{cls}L_{c}(\bm{z}^{\eta},c(\beta^{\eta}))\right)$
-where $\beta_{cls}$ is a tunable hyperparameter, and $L_{c}(\bm{z}^{\eta},c(\beta^{\eta}))$ is the corresponding loss term depending on the provided information.
-For instance, when taking $c$ as the classification label of $\beta^{\eta}$, $L_{c}$ can be taken as the cross-entropy loss.
-We note that the coefficient $\beta_{KL}$ is closely connected to the adjustable hyperparameter in $\beta-$VAE (Higgins et al., 2017).
-Taking a larger $\beta_{KL}$ encourages disentanglement (Locatello et al., 2019), while it also adds additional constraint on the implicit capacity of the latent bottleneck that results in information loss (Burgess et al., 2018).
+$L_{loss}=\dfrac{1}{S}\sum_{\eta=1}^{S}\Bigg(\beta_{d}\sum_{j=1}^{n_{train}}\left\|\hat{\mathcal{G}}[\bm{f}^{\eta}_{j};\hat{\theta}^{\eta}]-\bm{u}^{\eta}_{j}\right\|^{2}_{L^{2}({\Omega})}+\left\|\hat{\theta}^{\eta}-\mu^{\eta}_{\theta}\right\|^{2}+\beta_{KL}\left\|\mu^{\eta}_{\bm{z}}\right\|^{2}+\beta_{cls}L_{c}(\bm{z}^{\eta},c(\beta^{\eta}))\Bigg)$
+with $\beta_{cls}$ a tunable parameter and $L_{c}(\bm{z}^{\eta},c(\beta^{\eta}))$ the corresponding loss term based on provided information (e.g., $L_{c}$ is the cross-entropy loss when taking $c$ as the classification label of $\beta^{\eta}$).
+Note that $\beta_{KL}$ is closely connected to the adjustable hyperparameter in $\beta$-VAE (Higgins et al., 2017).
+Taking a larger $\beta_{KL}$ encourages disentanglement (Locatello et al., 2019), while it also adds additional constraint on the implicit capacity of the latent bottleneck resulting in information loss (Burgess et al., 2018).
 On the other hand, the data reconstruction term forces the latent factors to contribute to the “reconstruction” of the more complicated solution field in a global perspective, and hence increasing $\beta_{d}$ is anticipated to alleviate information loss.
-In our empirical studies, we demonstrate the interplay between these two tunable hyperparameters.
+In our empirical study, we demonstrate the interplay between these two tunable hyperparameters.
 4 Experiments
-We assess the performance of DisentangO across a variety of physics modeling and discovery datasets.
-In the following, we denote a DisentangO model with a latent dimension of $n$ as DisentangO-$n$.
-Our evaluation focuses on the following key aspects.
-Firstly, we demonstrate the capability of DisentangO in forward PDE learning, and compare its performance with a number of relevant baselines.
-In particular, we select three NO-related baselines, i.e., FNO (Li et al., 2020c), NIO (Molinaro et al., 2023) and MetaNO (Zhang et al., 2023), as well as two non-NO baselines, i.e., VAE (Kingma and Welling, 2013) and $\beta$-VAE (Higgins et al., 2017).
+Table 1: Test errors and the number of trainable parameters in experiment 1. Bold number highlights the best method.
+Models #param per-epoch Test errors
+time (s) data $\bm{z}$ (SC1)
+DisentangO 697k 12.2 1.65% 4.63%
+MetaNO 296k 9.8 1.59% -
+NIO 709k 5.6 - 15.16%
+FNO 698k 9.1 2.45% 14.55%
+UFNO 720k 21.2 7.61% 11.23%
+WNO 672k 183.5 8.09% 9.95%
+PIANO 699k 16.9 7.99% 15.35%
+CAMEL 654k 5.53 112.28% -
+SMDP 671k 5.12 - 17.76%
+InVAErt 707k 0.1 - 5.16%
+FUSE 706k 2.2 - 4.99%
+FUSE-f 707k 11.4 16.33% 6.19%
+VAE 698k 2.8 49.97% 16.34%
+convVAE 664k 2.8 81.11% 16.27%
+$\beta$-VAE 698k 2.8 51.16% 16.47%
+We assess DisentangO across various physics modeling and discovery datasets, denoting a model with latent dimension $n$ as DisentangO-$n$.
+Our evaluation focuses on several key aspects.
+Firstly, we demonstrate the capability of DisentangO in forward PDE learning, and compare its performance with a total of 14 relevant baselines.
+In particular, we select 8 NO baselines, i.e., FNO (Li et al., 2020c), UFNO (Wen et al., 2022), NIO (Molinaro et al., 2023), WNO (Tripura & Chakraborty, 2022), PIANO (Zhang et al., 2024), MetaNO (Zhang et al., 2023), FUSE (Lingsch et al., 2024) and its extension, as well as six non-NO baselines, i.e., CAMEL (Blanke & Lelarge, 2023), InVAErt (Tong et al., 2024), SMDP (Holzschuh et al., 2023), two VAE variants (Kingma & Welling, 2013) and $\beta$-VAE (Higgins et al., 2017).
 Secondly, we showcase the merits of DisentangO in inverse modeling for interpretable physics discovery.
-Thirdly, we perform parametric studies on the associated disentanglement hyperparameters.
-Lastly, we provide interpretations of the disentangled parameters and point out future directions.
-Further details are provided in Appendix C.
+We perform parametric studies on the associated disentanglement parameters.
+Lastly, we provide interpretations of the disentangled parameters.
+Further details and an additional experiment for identifiability demonstration are provided in Appendices B and C.4, respectively.
 4.1 Supervised Forward and Inverse PDE Learning
-Table 1: Test errors and the number of trainable parameters. Bold number highlights the best method.
-Models #param Test errors
-data latent supervision
-DisentangO 697,383 1.65% 4.63%
-MetaNO 295,842 1.59% -
-NIO 709,260 - 15.16%
-FNO 698,284 2.45% 14.55%
-VAE 697,672 49.97% 16.34%
-$\beta$-VAE 697,672 51.16% 16.47%
 We start by investigating DisentangO’s capability in solving both forward and inverse PDE problems in a fully supervised setting.
@@
 In this setting, the model takes as input the padded traction field and learns to predict the displacement.
-As the number of latent factors is fixed for supervised learning in this example, we defer our ablation study to the second experiment of semi-supervised learning.
-We report in Table 1 our experimental results from DisentangO as well as its comparison with the selected baselines.
-Note that only FNO, VAE and $\beta$-VAE are able to handle both forward and inverse problems, whereas MetaNO and NIO can only solve the forward problem and the inverse problem, respectively.
+As the number of latent factors is fixed for supervised learning in this experiment, we defer our ablation study to the second experiment.
+We report in Table 1 our experimental results.
+Note that only FNO, PIANO, FUSE-f, VAE and its variations are able to handle both forward and inverse problems, whereas MetaNO can only solve the forward problem, and NIO, InVAErt, and FUSE can only solve the inverse problem.
 With this caveat in mind, MetaNO achieves the best performance in forward PDE learning, with DisentangO performing comparably well and beating the third best model by 32.7% in accuracy.
-In terms of inverse modeling with full latent supervision, DisentangO is the only method that can hold the error below 5%, outperforming the second best method by 68.2%.
+In terms of inverse modeling with full latent supervision (SC1), DisentangO is the only method that can hold the error well below 5%, outperforming the second best method by 7.2% and the second best joint (i.e., simultaneous forward and inverse) solver by 25.2%.
 4.2 Semi-Supervised Mechanical MNIST Benchmark
-We next explore in the context of semi-supervised learning and apply DisentangO and relevant baselines to the benchmark Mechanical MNIST (MMNIST) dataset (Lejeune, 2020).
+We consider semi-supervised learning and apply DisentangO to the Mechanical MNIST (MMNIST) benchmark (Lejeune, 2020).
 MMNIST comprises 70,000 heterogeneous material specimens undergoing large deformation, each governed by a material model of the Neo-Hookean type with a varying modulus converted from the MNIST bitmap images.
@@
 Since only partial knowledge is available for each image (i.e., the corresponding digit), we apply semi-supervised learning to the latent factors to classify the digits.
+Table 2: Test errors and number of trainable parameters in experiment 2.
+DisentangO is abbreviated as DNO due to space limit.
+Bold number highlights the best method that can handle both forward and inverse settings.
+Models DNO-2 DNO-5 DNO-10 DNO-15 VAE $\beta$-VAE MetaNO PIANO
+#param (M) 0.66 0.97 1.49 2.02 2.02 2.02 0.35 2.05
+$\beta_{d}=1$ 12.82% 9.56% 7.36% 6.29% 16.34% 17.13% 2.68% 13.73%
+$\beta_{d}=10$ 11.51% 9.16% 6.62% 5.95% - - -
+$\beta_{d}=100$ 11.49% 8.43% 6.65% 5.48% - - -
+$\beta_{d}=1000$ 11.62% 8.22% 6.50% 5.80% - - -
 Ablation study.
-Firstly, we investigate DisentangO’s predictability in forward PDE learning by comparing its performance to MetaNO (i.e., the base meta-learnt NO model without the hierarchical structure).
+Firstly, we investigate DisentangO’s predictability in forward PDE learning by comparing its performance to MetaNO (i.e., the base meta-learned NO model without the hypernetwork structure).
 As seen in Table 2, MetaNO achieves a forward prediction error of 2.68%, which serves as the optimal bound for DisentangO.
@@
 In contrast, although the accuracy of DisentangO with latent semi-supervision slightly deteriorates due to the additional regularization effect from the classification loss term, it is able to correctly recognize the embedded digits and leverage this partial knowledge in disentangling meaningful latent factors.
-Table 2: Test errors and the number of trainable parameters in MMNIST benchmark.
-Models are abbreviated (0=DisentangO-2, 1=DisentangO-5, 2=DisentangO-10, 3=DisentangO-15, 4=VAE, 5=$\beta$-VAE, 6=MetaNO).
-Bold number highlights the best method that can handle both forward and inverse settings.
-Models 0 1 2 3 4 5 6
-#param (M) 0.66 0.97 1.49 2.02 2.02 2.02 0.35
-Error $\beta_{d}=1$ 12.82% 9.56% 7.36% 6.29% 16.34% 17.13% 2.68%
-$\beta_{d}=10$ 11.51% 9.16% 6.62% 5.95% - - -
-$\beta_{d}=100$ 11.49% 8.43% 6.65% 5.48% - - -
-$\beta_{d}=1000$ 11.62% 8.22% 6.50% 5.80% - - -
-Figure 2: MMNIST unsupervised scores against $\beta_{d}$ with DisentangO-2 (left) and DisentangO-15 (right).
-By comparing $\beta_{d}=1$ (solid lines) with $\beta_{d}=100$ (dashed lines), increasing $\beta_{d}$ forces the latent factors to maximize the contained information and in turn decreases MI, thus encouraging disentanglement.
-Figure 3: MMNIST scatterplot with DisentangO-2 and $\beta_{d}=1$: left: $(\beta_{kl}=1,\beta_{cls}=100)$ (data test error 18.81%), middle: $(\beta_{kl}=10,\beta_{cls}=10)$ (data test error 16.94%), right: fully unsupervised DisentangO-2 without classification loss $(\beta_{kl}=100)$ (data test error 12.65%).
 Comparison with additional baselines.
-We compare the performance of DisentangO with relevant additional baselines, i.e., VAE and $\beta$-VAE.
-We abandon the three NO-related baselines in this and the following experiments because they are black-box approximators and do not possess any mechanism to extract meaningful information without supervision.
-In this context, even the least accurate DisentangO-2 model (i.e., with $\beta_{d}=1$) with significantly fewer parameters outperforms the two selected baselines in Table 2 by 21.5% and 25.2%, respectively, with the best performing DisentangO-15 model beating the two baselines by 66.5% and 68.0% in accuracy, respectively.
+We compare DisentangO with additional baselines, i.e., VAE and $\beta$-VAE.
+We abandon the three NO-related baselines in this and the following experiments as they are black-box approximators and do not possess any mechanism to extract meaningful information without supervision.
+In this context, even the least accurate DisentangO-2 model (i.e., with $\beta_{d}=1$) with significantly fewer parameters outperforms the selected baselines in Table 2 by 21.5% and 25.2%, respectively, with the best performing DisentangO-15 model beating the baselines by 66.5% and 68.0% in accuracy, respectively.
 We do not proceed to compare the physics discovery capability between them as the baseline models are considerably inaccurate in forward prediction.
@@
 4.3 Unsupervised Heterogeneous Material Learning
-We now demonstrate the applicability of DisentangO in unsupervised learning in the context of learning synthetic human tissues that exhibit highly organized structures, with collagen fiber arrangements varying spatially.
-In this subject, it is critical to understand the underlying low-dimensional disentangled properties in the hidden latent space of this complex, high-dimensional microstructure, as inferred from experimental mechanical measurements.
+Table 3: Test errors and number of trainable parameters in experiment 3.
+DisentangO is abbreviated as DNO due to space limit.
+Bold number highlights the best method for both forward and inverse settings.
+Models #param (M) Test error
+$\beta_{d}=1$ $\beta_{d}=100$
+DNO-2 0.63 26.33% 25.18%
+DNO-5 0.94 17.51% 15.75%
+DNO-10 1.46 10.76% 10.01%
+DNO-15 1.98 7.11% 7.02%
+DNO-30 3.55 5.33% 5.28%
+VAE 3.55 61.10% -
+$\beta$-VAE 3.55 57.04% -
+MetaNO 0.32 2.67% -
+PIANO 3.42 50.71%
+We demonstrate DisentangO in unsupervised learning in the context of learning synthetic tissues that exhibit highly organized structures, with collagen fiber arrangements varying spatially.
+In this case, it is critical to understand the underlying low-dimensional disentangled properties in the hidden latent space of complex, high-dimensional microstructure, as inferred from experimental mechanical measurements.
 We generate two datasets, each containing 500 specimens and 100 loading/displacement pairs.
 The first dataset features variations in fiber orientation distributions using a Gaussian Random Field (GRF) and the second differs fiber angles in two segmented regions, separated by a centerline with a randomly rotated orientation.
-Figure 4: Latent traversal of DisentangO-3 in unsupervised heterogeneous material learning, where the three latent dimensions control the border rotation between the two segments (top), the relative fiber orientation between the two segments (middle), and the fiber orientation of the top segment (bottom), respectively.
-Legend indicates fiber orientation ranging from 0 to $\pi$.
-Table 3: Test errors and the number of trainable parameters in experiment 3.
-Models are abbreviated (0=DisentangO-2, 1=DisentangO-5, 2=DisentangO-10, 3=DisentangO-15, 4=DisentangO-30, 5=VAE, 6=$\beta$-VAE, 7=MetaNO).
-Bold number highlights the best method that can handle both forward and inverse settings.
-Models #param (M) Error
-$\beta_{d}=1$ $\beta_{d}=100$
-0 0.63 26.33% 25.18%
-1 0.94 17.51% 15.75%
-2 1.46 10.76% 10.01%
-3 1.98 7.11% 7.02%
-4 3.55 5.33% 5.28%
-5 3.55 61.10% -
-6 3.55 57.04% -
-7 0.32 2.67% -
-Figure 5: Unsupervised MI score against $\beta_{kl}$ with DisentangO-2 in heterogeneous material learning.
 We report in Table 3 our experimental results of DisentangO with different latent dimensions and data loss strength $\beta_{d}$, along with comparisons with the baseline model results.
@@
 5 Conclusion
-We present DisentangO, a novel hyper-neural operator architecture designed to disentangle latent physical factors embedded within black-box neural operator parameters.
-DisentangO leverages a hierarchical neural operator architecture that extracts varying parameters of the governing PDE through a task-wise adaptive layer, then further disentangles these variations into distinct latent factors.
-By learning these disentangled representations, DisentangO not only enhances physical interpretability but also enables more robust generalization across diverse physical systems.
-We perform empirical evaluations across supervised, semi-supervised, and unsupervised learning contexts.
-Limitations: Due to limited computational resource, our experiments are focused on learning a small-to-medium number ($<500$).
-Expanding the coverage to encompass a broader range of physical systems and enabling learning across more diverse parameters would be beneficial.
-Acknowledgments and Disclosure of Funding
+We present DisentangO for disentangling latent physical factors embedded within black-box NO parameters.
+DisentangO leverages a hypernetwork-type NO architecture that extracts varying parameters of the governing PDE through a task-wise adaptive layer, and further disentangles these variations into distinct latent factors.
+By learning disentangled representations, DisentangO not only enhances physical interpretability but also enables robust generalization across diverse physical systems under different learning contexts.
+Limitations: Since the scalability of DisentangO is governed by the scalability of the NO backbone used for forward modeling, which has been extensively discussed in previous works, this study focuses on experiments involving high latent dimensions.
+Demonstrations on high-dimensional PDEs are beyond the scope of the current work.
+Acknowledgments
 L.
-Zhang would like to acknowledge support by the AFOSR grant FA9550-22-1-0197.
+Zhang was partially funded by the NSF grant DMS 2436624.
 Y.
-Yu would like to acknowledge the support by the National Science Foundation under award DMS-2436624 and the support from the National Institute of Health under award 1R01GM157589-01.
-Portions of this research were conducted on Lehigh University’s Research Computing infrastructure partially supported by NSF Award 2019035.
-Appendix A Proofs of identifiability analysis
+Yu would like to acknowledge the support by the National Institute of Health under award 1R01GM157589-01.
+Portions of this research were conducted on Lehigh University’s Research Computing infrastructure, partially supported by NSF Award 2019035.
+Appendix A Identifiability analysis
+A.1 Proof of the main theorems
 We first provide the proof of Theorem 1:
@@
 Denoting $r:={\mathcal{G}}_{f}\circ\hat{\mathcal{G}}^{-1}_{f}$, it is the transformation between the true $\theta$ and the estimated one, and it is invertible and invariant with respect to $\bm{f}$.
-We proceed to derive the relation between $\bm{z}$ and $\hat{\bm{z}}$: since $\theta=r(\hat{\theta})$, with the invertibility assumption $\theta=\mathcal{H}^{-1}(\bm{z})$ and $\hat{\theta}=\hat{\mathcal{H}}^{-1}(\hat{\bm{z}})$, we obtain:
-$\bm{z}=\mathcal{H}(\theta)=\mathcal{H}\circ r(\hat{\theta})=\mathcal{H}\circ r\circ\hat{\mathcal{H}}^{-1}(\hat{\bm{z}}).$
+[…]
+[…]
 Denoting $h:=\mathcal{H}\circ r\circ\hat{\mathcal{H}}^{-1}$, it is the transformation between the true latent variable and the estimated one, and it is invertible because $r$, $\mathcal{H}$ and $\hat{\mathcal{H}}$ are all invertible.∎
@@
 Differentiating the above equation twice with respect to $\hat{z}_{k}$ and $\hat{z}_{q}$, $k\neq q$, yields
-$\sum_{i}\left(\dfrac{\partial^{2}q_{i}(z_{i},\bm{b})}{\partial z_{i}^{2}}\dfrac{\partial z_{i}}{\partial\hat{z}_{k}}\dfrac{\partial z_{i}}{\partial\hat{z}_{q}}+\dfrac{\partial q_{i}(z_{i},\bm{b})}{\partial z_{i}}\dfrac{\partial^{2}z_{i}}{\partial\hat{z}_{k}\partial\hat{z}_{q}}\right)+\dfrac{\partial^{2}\log|J_{h}|}{\partial\hat{z}_{k}\partial\hat{z}_{q}}=0\text{ .}$ (A.1)
+[…]
 To show the identifiability, one can rewrite the Jacobian $J_{h}$ as:
@@
 Taking $\bm{b}=\bm{b}^{0},\cdots,\bm{b}^{2d_{z}}$ in equation A.1 and subtracting them from each other, we have
-$\sum_{i=1}^{d_{z}}\left(\left(\dfrac{\partial^{2}q_{i}(z_{i},\bm{b}^{j})}{\partial z_{i}^{2}}-\dfrac{\partial^{2}q_{i}(z_{i},\bm{b}^{0})}{\partial z_{i}^{2}}\right)\dfrac{\partial z_{i}}{\partial\hat{z}_{k}}\dfrac{\partial z_{i}}{\partial\hat{z}_{q}}+\left(\dfrac{\partial q_{i}(z_{i},\bm{b}^{j})}{\partial z_{i}}-\dfrac{\partial q_{i}(z_{i},\bm{b}^{0})}{\partial z_{i}}\right)\dfrac{\partial^{2}z_{i}}{\partial\hat{z}_{k}\partial\hat{z}_{q}}\right)=0\text{ ,}$
+[…]
 where $j=1,\cdots,2d_{z}$. With the linear independence condition for $w$, this is a $2d_{z}\times 2d_{z}$ linear system, and therefore the only solution is
@@
 ∎
+A.2 Further discussion on the assumptions
+Herein, we provide additional discussion on the validity and emprical validation for Assumptions 1-4.
+As seen in the proof above, Assumptions 1 (Density Smoothness and Positivity) and 2 (Invertibility) are required to guarantee that there exists a smooth and injective mapping $h:=\mathcal{H}\circ r\circ\hat{\mathcal{H}}^{-1}$, from the ground-truth latent embedding $\bm{z}$ to the learned embedding $\hat{\bm{z}}$.
+Furthermore, the smoothness assumption further makes it feasible to take derivatives of $\bm{z}$ with respect to $\hat{\bm{z}}$, which supports the permutation-wise identifiability proof for Theorem 2.
+Here, we note that the smoothness assumption may possibly be relaxed to $C^{2}$.
+Assumptions 3 (Conditional Independence) and 4 (Linear Independence) are needed to show that the Jacobian of $h$ has one and only one non-zero component for each column.
+Without these assumptions, it is possible that the data from different $\mathbf{b}$ lack variability.
+While the first three assumptions are common in many VAE architectures, the last assumption is plausible for many real-world data distributions.
+For instance, when the prior on the latent variables $p(\bm{z}|\bm{b})$ is conditionally factorial, where each element $z_{i}$ has a univariate exponential family distribution given conditioning variable $\bm{b}$:
+$p(\bm{z}|\bm{b})=\prod_{i}\dfrac{Q_{i}(z_{i})}{Z_{i}(\bm{b})}\exp\left[\sum_{j=1}^{k}T_{i,j}(z_{i})\lambda_{i,j}(\bm{b})\right],$
+where $Q_{i}$ is the base measure, $Z_{i}(\bm{b})$ is the normalizing constant, $T_{i,j}$ are the sufficient statistics, and $\lambda_{i,j}$ are the corresponding parameters depending on $\bm{b}$.
+This exponential family has universal approximation capabilities.
+Additionally, we note that this distribution is conditionally independent with
+$q_{i}=\log(Q_{i}(z_{i}))-\log(Z_{i}(\bm{b}))+\left[\sum_{j=1}^{k}T_{i,j}(z_{i})\lambda_{i,j}(\bm{b})\right],$
+and the linear independence indicates that the matrix formed by
+$\omega(\bm{z},\bm{b}^{\eta})-\omega(\bm{z},\bm{b}^{0})=\left(\mathbf{T}^{\prime}(\lambda(\bm{b}^{\eta})-\lambda(\bm{b}^{0})),\mathbf{T}^{\prime\prime}(\lambda(\bm{b}^{\eta})-\lambda(\bm{b}^{0}))\right),\,\eta=1,\cdots,S$
+has full rank $2d_{z}$.
+In the fully supervised case, the conditional independence and linear independence assumptions are automatically guaranteed by picking proper distributions $p(\bm{z}|\bm{b})$ when designing the VAE architecture.
+In the semi-supervised and unsupervised cases, one can also validate these assumptions when the true values of $\bm{b}$ is given on some tasks, by inferring an empirical distribution of $p(\bm{z}|\bm{b})$ from these tasks.
+To investigate such a capability, in the additional synthetic experiment in Appendix C.4, we consider an unsupervised setting, estimate $p(\bm{z}|\bm{b})$ from the trained model, and check the linear independence condition by calculating the vector in equation 3.15 for each task $\bm{b}^{\eta}$ and forming an $S\times 2d_{z}$ matrix from all tasks.
+When the rank of this matrix is $2d_{z}$, it means that we can select $2d_{z}+1$ tasks from them with sufficient variability, such that the linear independence condition is satisfied.
+As a demonstration, in Appendix C.4 we validate this assumption on a synthetic dataset and show that the identifiability can be largely achieved with Gaussian distributions of distinct means and variances.
 Appendix B Additional experimental details, results and discussion
-Herein, we provide additional details in training and baseline models, as well as more results as a supplement of Section 4.
+We provide additional details in training and baseline models, as well as more results as a supplement of Section 4.
+Figure 3: MMNIST unsupervised scores against $\beta_{d}$ with DisentangO-2 (left) and DisentangO-15 (right).
+By comparing $\beta_{d}=1$ (solid lines) with $\beta_{d}=100$ (dashed lines), increasing $\beta_{d}$ forces the latent factors to maximize the contained information and in turn decreases MI, thus encouraging disentanglement.
+Figure 4: Unsupervised MI score against $\beta_{kl}$ with DisentangO-2 in heterogeneous material learning.
+Figure 5: Latent traversal of DisentangO-3 in unsupervised heterogeneous material learning, where the three latent dimensions control the border rotation between the two segments (top), the relative fiber orientation between the two segments (middle), and the fiber orientation of the top segment (bottom), respectively.
+Legend indicates fiber orientation ranging from 0 to $\pi$.
 B.1 Training details
-In all experiments, we adopt the Adam optimizer for optimization and use a 16-layer DisentangO model.
+In all experiments, we adopt the Adam optimizer for optimization and use a $n$-layer DisentangO model, where $n=8$ in the first experiment and $n=16$ in the second and third experiments due to increased complexity.
 For fair comparison across different models, we tune the hyperparameters, including the learning rates, the decay rates, and the regularization parameters, to minimize the validation loss.
@@
 1:Denote data reconstruction loss $L_{data}$, task-wise NO parameter reconstruction loss $L_{recon}$, KL loss $L_{KL}$ and semi-supervision loss $L_{semi}$ as:
-$L_{data}=$ $\dfrac{1}{S}\sum_{\eta=1}^{S}\left[-E(\log(p(\bm{u}|\bm{f},\theta^{\eta})))\right]\text{ ,}\quad L_{recon}=\dfrac{1}{S}\sum_{\eta=1}^{S}\left[-E_{q(\bm{z}^{\eta}|\theta^{\eta})}\log p(\theta^{\eta}|\bm{z}^{\eta})\right]\text{ ,}$
-$L_{KL}=$ $\dfrac{1}{S}\sum_{\eta=1}^{S}\left[D_{KL}\left(q(\bm{z}^{\eta}|\theta^{\eta})||p(\bm{z}^{\eta})\right)\right]\text{ ,}\quad L_{semi}=L_{c}(\bm{z}^{\eta},c(\bm{b}^{\eta}))\text{ .}$
-2:SC1: Unsupervised/SC3: Supervised
+[…]
+[…]
+2:SC1: Supervised/SC3: Unsupervised
 3:The total loss is comprised of the data loss, the NO parameter reconstruction loss, and the KL loss: $L_{loss}=\beta_{d}L_{data}+L_{recon}+\beta_{KL}L_{KL}.$
@@
 The parameter of each baseline is given in the following, where the parameter choice of each model is selected by tuning the number of layers and the width (channel dimension), keeping the total number of parameters on the same magnitude.
-  • MetaNO: We use a 16-layer IFNO model with the lifting layer as the adaptive layer.
+  • MetaNO: We use a 8-layer IFNO model with the lifting layer as the adaptive layer.
 We keep the total number of parameters in MetaNO the same as the number of parameters used in forward PDE learning in DisentangO.
   • NIO: We closely follow the setup in Molinaro et al.
-(2023) and use four convolution blocks as the encoder for the branch net and a fully connected neural network with two hidden layers of 256 neurons as the trunk net, with the number of basis functions set to 50.
+(2023), where two neural operators (DeepONet and FNO) are stacked together to realize the operator-to-function intuition.
+The first operator maps multiple solution functions to a set of representations (which can be seen as an analog of eigenfunctions), and the second operator infers the underlying parameter field from the mixed representations.
+As NIO requires the solution field as input, it cannot be used as a forward solver.
+Hence, NIO only solves the inverse PDE problem, and it can only be applied to the fully supervised setting.
+Specifically, we use four convolution blocks as the encoder for the branch net and a fully connected neural network with two hidden layers of 256 neurons as the trunk net, with the number of basis functions set to 50.
 For the FNO part, we use one Fourier layer with width 32 and modes 8, as suggested in Molinaro et al.
 (2023).
-  • FNO: We adopt a 4-layer FNO with width 26 and modes 8.
-For the forward model, in addition to the loading field and coordinates as input, we also concatenate the ground-truth material properties to form the final input.
+  • FNO: Since FNO is originally designed as a function-to-function mapping, we consider the inverse optimization procedure following Lee et al.
+(2024), and develop a two-phase process to solve the forward and inverse problems sequentially.
+In the first phase, we construct the forward mapping from the loading field $\bm{f}$ and the ground-truth material parameter $\bm{b}$ to the corresponding solution $\bm{u}$ as: $\mathcal{G}^{FNO}[\bm{f},\bm{b};\theta^{FNO}](\bm{x})=\bm{u}(\bm{x})$.
+This can be seen as an analog of the forward solution operator $\mathcal{G}$ in our setting.
+Then, with the trained FNO as a surrogate for the forward solution operator, we fix its NN parameters $\theta^{FNO}$, and use it together with gradient-based optimization to solve for the optimal material parameters as an inverse solver.
+Specifically, given a set of loading/solution data pairs $\{(f_{i},u_{i})\}_{i=1}^{N}$, we start from a random guess of the underlying material parameters (typically chosen as the average of all available instances of material parameters for fast convergence), and minimize the difference between the predicted displacement field from FNO and the ground-truth one:
+    $b^{*}=\text{argmin}_{b}\sum_{i=1}^{N}\|u_{i}-\mathcal{G}^{FNO}[f_{i},b;\theta^{FNO}]\|^{2}.$
+    As the FNO parameters are fixed, we can back propagate this loss and optimize the input material parameters in an iterative fashion.
+We adopt a 4-layer FNO with width 26 and modes 8.
+For the forward model, in addition to the loading field and the coordinates as input, we also concatenate the ground-truth material properties to form the final input.
 For the inverse model, we employ an iterative gradient-based optimization to solve for the optimal material parameters for physics discovery.
+For fair comparison, in terms of the averaged per-epoch runtime, we report the sum of both the forward and inverse solvers.
+The averaged per-epoch runtimes for the forward solver and the inverse solver are 6.2 seconds and 2.9 seconds, respectively, accounting for the total per-epoch runtime of 9.1 seconds in Table 1.
+  • UFNO: The 2D U-FNO model extends the Fourier Neural Operator (FNO) architecture by incorporating U-Net-style skip connections.
+The network consists of six spectral convolution blocks, each combining a global Fourier operator and a local $1\times 1$ convolution.
+In the later three layers, additional skip-enhanced feature extraction is provided by U-Net blocks.
+The input consists of spatial features $f(x,y)$, concatenated with coordinate embeddings $(x,y)$ and 5 material parameters.
+The input is lifted to a higher-dimensional latent space via a linear layer of size $[8,20]$.
+Each of the six layers performs the following composite operation:
+    $v_{j+1}=\text{ReLU}(\mathcal{F}_{j}(v_{j})+\mathcal{W}_{j}(v_{j})+\mathcal{U}_{j}(v_{j})),$
+    where $\mathcal{F}_{j}$ is a spectral convolution layer (2D Fourier transform, mode truncation, linear transformation, and inverse transform), $\mathcal{W}_{j}$ is a $1\times 1$ convolution, and $\mathcal{U}_{j}$ is a U-Net block (included only in layers 4 to 6).
+After all six layers, the output is projected through two fully connected layers of sizes $[20,128]$ and $[128,2]$ to produce the final 2-channel output.
+  • WNO: The Wavelet Neural Operator (WNO) consists of 4 stacked wavelet kernel integral layers.
+The input is first lifted to a high-dimensional representation using a fully connected layer of size $[8,27]$.
+Each of the 4 layers combines a learned local operator $W$ (implemented as $1\times 1$ convolution) with a non-local wavelet-based integral operator $K$ (via continuous wavelet transform).
+The architecture follows: The input $f(x,y)$ is first augmented by concatenating it with the positional grid coordinates $(x,y)$ and additional 5 material parameters.
+This augmented input is then lifted into a higher-dimensional space via a fully connected layer that maps the input to a hidden width of width.
+The core of the architecture consists of 4 wavelet-based layers, the output is projected back to the desired output space using two linear layers of sizes $[27,128]$ and [128,41412][128,41*41*2].
+Mish activations are applied after each intermediate layer except the last projection layer.
+  • CAMEL: The architecture consists of two networks, V-Net and C-Net
+    $G(f(x);\theta,w)=c((f(x);\theta)+w^{\top}v(f(x);\theta).$ (B.1)
+    The score network $V_{\phi}$ is a 5-layer MLP with width 320 and Tanh activations, with layer sizes $[d,320,320,320,320,r]$ The coefficient network $c_{\theta}$ is a 3-layer MLP with width 128 and Tanh activations, with layer sizes $[d,128,128,1]$,, where $d$ is the size of the flatten $f(x)$, and $r=5$
+  • SMDP: Since we do not have an explicit physical operator to evaluate $P^{-1}(z,f(x))$, we use only the score network, implemented as a 6-layer MLP with layer sizes $128,512,256,128,64,32,d_{z}]$, where $d_{z}$ denotes the dimension of the latent space.
+  • PIANO: We use the ground-truth material properties as the physical-invariant embedding.
+The loading field is first processed by a lifting layer, while the material parameters are passed through an attention module implemented as a 3-layer MLP with hidden dimension 32.
+The outputs of the lifting layer and the attention module are then concatenated and fed into a convolutional layer with width 26, followed by a 3-layer IFNO with width 26 and modes 8.
+For both the forward and inverse solution procedures, we adopt the same settings as in the FNO baseline.
+  • InVAErt: We directly take the InVAErt implementation from Lingsch et al.
+(2024) and define the encoder, the VAE encoder and the decoder as 4-layer MLPs with hidden dimension 96 and silu activation function.
+  • FUSE: For the forward model, we take three Fourier layers in addition to the first band-limited lifting layer that increases the dimension of the parameters and performs an inverse Fourier transform.
+On the other hand, the inverse model maps the functional input to the parameter space by employing a concatenation of two Fourier layers and a band-limited forward Fourier transform that generates a fixed-size latent representation of the input function, followed by a flow-matching posterior estimation (FMPE) that maps to the final parameter output.
+All Fourier layers have a latent width of 32 and 8 modes, while the FMPE flow has 4 layers of width 360.
+  • FUSE-f: Since the original FUSE model assumes a constant loading field, it cannot handle situations where the input loading field changes that create multiple instances of a PDE system.
+We therefore create a FUSE variation (denoted as “FUSE-f”) that takes a concatenation of both the displacement field and the loading field as input to the inverse model.
+For the forward model, we concatenate the loading field to the output of the first band-limited lifting layer and subsequently use a one-layer MLP to map it back to the original dimension.
+All other settings are the same as the original FUSE baseline.
   • VAE: We use a 2-layer MLP of size [1681, 136, 30] as the encoder, and another 2-layer MLP of size [30, 136, 1681] as the decoder, with the size of the bottleneck layer being 30.
+  • convVAE: We use a convolutional layer with 136 kernels of size $3\times 3$ with a stride of 2 pixels and a fully connected layer of size [59976, 5] as the encoder, and a fully connected layer of size of [5, 59976] and a transposed convolutional layer with 2 kernels of size $3\times 3$ as the decoder.
   • $\beta$-VAE: The parameter choice of the $\beta$-VAE baseline is the same as the VAE baseline, except that we tune the $\beta$ hyperparameter.
+Model-agnostic architecture.
+The proposed DisentangO architecture is model-agnostic and can incorporate any neural operator as the forward solver, including FNO, UFNO, and WNO.
+This flexibility stems from DisentangO’s design, which only requires designating the lifting layer as the task-wise adaptation layer and adding the VAE structure for latent disentanglement (i.e., the encoder and the first decoder in Figure 1).
+DisentangO’s overall performance scales directly with the underlying neural operator’s capabilities.
+As long as the chosen NO achieves comparable forward prediction accuracy to IFNO, DisentangO will maintain similar performance, since (1) the upper bound of forward prediction accuracy depends on the base neural operator, and (2) the inverse prediction accuracy depends on how well the latent factors are encoded in the task-wise lifting layer parameters.
+We select IFNO for three key advantages: (1) parameter efficiency: the layer-independent parameter setting significantly reduces trainable parameters compared to alternatives, (2) theoretical guarantees: IFNO is a provably universal approximator as PDE solution operators, and (3) natural integration: seamless compatibility with MetaNO’s lifting layer approach for task-wise adaptation.
+If the IFNO part is replaced with other neural operators such as the considered baselines of FNO, UFNO, and WNO as in Table 1, the forward prediction accuracy will decrease (cf.
+the data test errors in Table 1), meaning that the upper bound of DisentangO’s forward prediction accuracy will decrease.
+Additionally, the inverse prediction accuracy (i.e., the z test error in Table 1) will likely drop as well, because the degraded forward prediction accuracy typically indicates degraded encoding of latent information in the lifting layer, which will negatively impact the inverse modeling even if the VAE part can reconstruct the lifting layer parameters perfectly.
+As IFNO performs the best in forward prediction and requires the least amount of trainable parameters, we choose IFNO as the forward prediction backbone in DisentangO.
+To demonstrate DisentangO’s compatibility with alternative neural operators, we replace IFNO with UFNO in the DisentangO architecture. The performance comparison is presented in Table 4.
+Table 4: Test errors and the number of trainable parameters in experiment 1. Bold number highlights the best method.
+Models #param per-epoch Test errors
+time (s) data $\bm{z}$ (SC1)
+DisentangO 697k 12.2 1.65% 4.63%
+MetaNO 296k 9.8 1.59% -
+MetaUFNO 304k 11.4 3.55% -
+DisentangO-UFNO 688k 14.2 6.61% 9.85%
+UFNO 720k 21.2 7.61% 11.23%
 B.3 Mechanical MNIST benchmark
 Figure 6: MMNIST latent traversal based on a randomly picked loading field as DisentangO-2 input.
-Herein, we demonstrate the latent traversal based on a randomly picked loading field as DisentangO-2 input.
+We demonstrate the latent traversal based on a randomly picked loading field as DisentangO-2 input.
 One can clearly see that the digit changes from “6” to “0” and then “2” from the top left moving down, and from “6” to “1” and then “7” moving to the right.
@@
 We also provide two exemplary MMNIST latent interpretations of the learned DisentangO-15 in Figures 7 and 8.
-Figure 7: Exemplary MMNIST latent interpretation of DisentangO-15: dimension 10 encodes the information for digit ‘8’, as is evidenced by the fact that all $y<-0.4$ regions on the scatterplots contain only digit ‘8’.
-Figure 8: Exemplary MMNIST latent interpretation of DisentangO-15: dimension 12 encodes the information for digit ‘3’, as is evidenced by the fact that all $y<-0.45$ regions on the scatterplots contain only digit ‘3’.
+Figure 7: Exemplary MMNIST latent interpretation of DisentangO-15: dimension 10 encodes the information for digit ‘8’, as is evidenced by the fact that all y<0.4y<-0.4 regions on the scatterplots contain only digit ‘8’.
+Figure 8: Exemplary MMNIST latent interpretation of DisentangO-15: dimension 12 encodes the information for digit ‘3’, as is evidenced by the fact that all y<0.45y<-0.45 regions on the scatterplots contain only digit ‘3’.
+B.4 Latent visualization and physical interpretation
+As the hidden parameter field $b^{\eta}$ of the PDE is generally not accessible, especially in the semi-supervised or unsupervised settings, one cannot directly reconstruct $b^{\eta}$ from the learned latent variables $z$.
+Through disentangled representation learning, the latent variables $z$ are anticipated to contain critical information of $b^{\eta}$, but there does not exist a direct and explicit mapping from $z$ to $b^{\eta}$.
+As a result, one cannot directly reconstruct $b^{\eta}$ from $z$.
+However, there are several tricks one can play with to obtain meaningful interpretations from the learned $z$.
+On one hand, one can feed a randomly selected input into DisentangO and manually define a desired $z$ in the latent space and perform a forward pass using the trained DisentangO.
+$b^{\eta}$ can then be visualized from the output, as demonstrated in Figure 6 in the MMNIST experiment.
+On the other hand, one can also train a simple MLP and construct a mapping from $z$ to $b^{\eta}$, provided that $b^{\eta}$ is available.
+With this mapping at hand, one can then traverse $z$ and visualize what each dimension of $z$ controls.
+This is demonstrated in Figure 5 in the third experiment.
+B.5 Further discussion
+Probabilistic latent space. We use a variational (probabilistic) encoder rather than a deterministic map for two reasons:
+  • Regularization and stability of the inverse map.
+The inverse operator in ill-posed PDE settings is highly sensitive: small perturbations in the solution $u$ can correspond to large changes in the inferred parameters.
+The VAE’s KL term imposes a distributional prior on the latent variables, preventing the encoder from collapsing to arbitrarily sharp or unstable mappings.
+This yields a smoother, more stable inverse that generalizes better under noise and limited data.
+  • Identifiability and disentanglement.
+The identifiability results we reference (e.g., up to component-wise invertible transformations) rely on the latent distribution having a simple, factorized prior.
+The VAE provides exactly this structure: a normalized, independent latent prior that constrains the representation and supports the theoretical guarantees.
+Scalability.
+Scaling interpretable neural operators to complex, high-dimensional PDEs is critical for broader impact.
+Our work is structured to first establish the foundational capabilities of DisentangO in controlled settings where disentanglement and representation quality can be rigorously verified.
+Notably, our work tackles scenarios with partially or fully unknown physics: the second experiment demonstrates DisentangO’s ability to disentangle and identify latent codes when only partial knowledge is available, and the third experiment aims to mimic a real-world fully unsupervised scenario where the physical parameters are entirely unknown, yet DisentangO successfully extracts interpretable latent factors.
+DisentangO is designed to be inherently scalable for several key reasons: (1) Modular architecture: DisentangO inherits the computational efficiency of the underlying neural operator backbone (IFNO in our implementation), which supports batched, GPU-accelerated inference in high dimensions.
+(2) Constant latent dimension: the latent code dimensionality remains low regardless of spatial resolution or system complexity, enabling practical inverse inference even for complex systems.
+(3) Linear scaling: the latent inference network scales linearly with the number of PDE instances, independent of spatial dimensionality.
+(4) Plug-and-play design: DisentangO can incorporate any state-of-the-art neural operator as its backbone, inheriting its scalability properties.
+In essence, the scalability of DisentangO is governed by the scalability of the neural operator used for forward modeling.
+As long as the NO component performs well and the training data spans sufficient variability in physical system parameters, DisentangO remains applicable.
+Encoding $b$ in lifting layers.
+Our choice to encode the parameters $b$ in the lifting layers $\theta_{P}$ is justified through three complementary perspectives.
+(1) Theoretical expressivity: as demonstrated in MetaNO (Zhang et al., 2023), varying only the lifting layer across tasks is sufficient to universally approximate a wide class of parametric operators.
+DisentangO inherits this universal approximation property while maintaining a meta-learning framework where the core iterative Fourier layers serve as a shared meta-operator across tasks.
+(2) Representation identifiability: restricting adaptation to the lifting layer significantly enhances the identifiability of latent representations, a critical requirement for disentangled inverse learning.
+When adaptation is allowed in deeper layers, the mapping between latent codes and operators becomes more diffuse and less interpretable.
+(3) Empirical validation: our experiments across diverse PDE types confirm that this design choice enables DisentangO to: (a) accurately model task-specific operators, (b) recover interpretable physical factors, and (c) generalize to unseen parameter combinations and interpolation tasks.
+Interpretability beyond parameter reconstruction.
+Our approach provides interpretability at the level of identifiable physical factors, which is often more meaningful than raw parameter reconstruction.
+Many physical parameters are inherently non-identifiable from observational data alone (e.g., material properties in regions with zero stress).
+Our method identifies the factors that actually govern system behaviors.
+Moreover, our experiments demonstrate clear interpretable patterns: (1) MMNIST (Figure 6): latent traversal reveals smooth transitions between digit patterns, showing the method captures meaningful morphological variations, (2) Heterogeneous materials (Figure 5): each latent dimension controls specific physical aspects: border rotation, relative fiber orientation, and segment-wise fiber orientation, (3) Semi-supervised results (Figure 2): latent factors successfully cluster different digit classes without explicit supervision.
+Our theoretical results on component-wise identifiability (Theorems 1 and 2) guarantee that the learned factors $\hat{z}$ correspond to true generative factors $z$ up to invertible transformations.
+This provides theoretical grounding for interpretability claims.
+In terms of practical utility, even without exact $b$ reconstruction, the disentangled factors can enable system classification and similarity assessment, design space exploration through latent traversal, transfer learning to new parameter regimes, and anomaly detection in physical systems.
+Quantitative interpretability metric.
+The evaluation of interpretability depends critically on the availability of ground truth, which varies across our three practical scenarios, i.e., supervised, semi-supervised, and unsupervised.
+In the fully supervised setting corresponding to experiment 1, When true generative factors are available, we use quantitative metrics.
+The interpretability metric is the latent supervision test error (cf.
+Table 1), which directly measures how well the learned latent variables $z$ recover the true physical parameters $b$.
+In the semi-supervised setting corresponding to experiment 2, the goal of interpreting physics is to discover the underlying microstructure governing the deformation.
+Under this setting, since the true generative factors are not available, one cannot use any closed-form metric to evaluate interpretability.
+However, since the ground-truth microstructure generation is controlled by MNIST digits, successful disentanglement should reveal this digit information.
+We thus follow the standard evaluation methods in disentangled representation learning and investigate if the learned latent variables indeed contain the digit information via latent traversal, and Figure 6 confirms that our learned latent variables capture the underlying digit structure, demonstrating physical interpretability through microstructure discovery.
+In the third setting of unsupervised learning, without ground truth labels, we evaluate interpretability through latent traversal analysis.
+Figure 5 demonstrates that our method discovers meaningful physical factors: border rotation between segments, relative fiber orientation between segments, and fiber orientation of individual segments—all physically meaningful properties for understanding the microstructure mechanism.
+Physical insights across different PDE types.
+Our method is designed for parametric PDE families—scenarios where underlying physics follows the same governing equations with varying parameters.
+This represents a very common and scientifically important class of problems in computational physics and engineering.
+While our current framework focuses on parametric variation within PDE families, the core architectural principles can extend to different PDE types.
+For example, in multi-physics systems, our approach could handle different but related physics (e.g., heat conduction vs.
+diffusion-reaction) by learning shared latent factors capturing common physical principles while using task-specific decoders.
+Extending to heterogeneous PDE types represents an exciting research frontier that would require modular architectures for different equation types, shared representation learning across physics domains, and physics-informed constraints for cross-domain transfer.
+Accuracy-interpretability trade-off.
+As discussed, MetaNO represents the forward prediction upper-bound for Disentango.
+The accuracy gap between the two is primarily due to insufficient latent dimensions relative to data complexity.
+Taking experiment 3 for instance, the data generation process may involve more than 30 true generative factors due to numerical solver noise and system complexity.
+Table 3 clearly shows the error decreases from 25.18% to 5.28% as we increase latent dimensions from 2 to 30—a trend that would continue with higher dimensions.
+We stopped at 30 dimensions because: (1) 5.28% error is reasonably accurate for practical use, and (2) limited computational resource.
+In practice, users should gradually increase latent dimensions until convergence to MetaNO’s performance, ensuring the model capacity matches the true generative factors.
+DisentangO prioritizes joint modeling with interpretability over pure forward accuracy.
+For applications requiring only forward prediction, specialized NOs remain more accurate.
+However, for scientific discovery where understanding mechanisms is essential, the slightly degraded accuracy with full interpretability represents a favorable trade-off.
 Appendix C Data Generation
-C.1 Example 1 - Supervised forward and inverse PDE learning
+C.1 Experiment 1 - Supervised forward and inverse PDE learning
 we consider the constitutive modeling of anisotropic fiber-reinforced hyperelastic materials governed by the Holzapfel–Gasser–Ogden (HGO) model, whose strain energy density function can be written as:
 $\eta$ $=\frac{E}{4(1+\nu)}(\bar{I}_{1}-2)-\frac{E}{2(1+\nu)}\ln J$
-$+\frac{k_{1}}{2k_{2}}(e^{k_{2}\langle S(\alpha)\rangle^{2}}+e^{k_{2}\langle S(-\alpha)\rangle^{2}}-2)+\frac{E}{6(1-2\nu)}(\frac{J^{2}-1}{2}-\ln J)\text{ ,}$ (C.1)
 […]
+where $\langle\cdot\rangle$ indicates the Macaulay bracket, $\alpha$, $k_{1}$ and $k_{2}$ are the fiber angle, modulus and exponential coefficient, respectively, $E$ denotes the Young’s modulus of the matrix, $\nu$ is Poisson’s ratio, and $S(\alpha)$ describes the fiber strain of the two fiber groups, $S(\alpha)=\frac{\bar{I}_{4}(\alpha)-1+|\bar{I}_{4}(\alpha)-1|}{2}$.
+$\bar{I}_{i}$ is the $i^{\text{th}}$ invariant of the right Cauchy-Green tensor $\bm{C}$, $\bar{I}_{1}=tr(\bm{C})$ and $\bar{I}_{4}=\bm{n}^{T}(\alpha)\bm{C}\bm{n}(\alpha)$, with $\bm{n}(\alpha)=[\cos(\alpha),\sin(\alpha)]^{T}$.
+In this context, the data generation process is controlled by sampling the material set $\{E,\nu,k_{1},k_{2},\alpha\}$, and the latent factors can be learned consequently in a supervised fashion.
+The physical parameters are sampled from $\frac{k_{1}}{k_{2}}\sim\mathcal{U}[90,100]$ $k_{2}\sim\mathcal{U}[0.001,0.1]$, $E\sim\mathcal{U}[0.5001,0.6001]$, $\nu\sim\mathcal{U}[0.2,0.3]$, and $\alpha\sim\mathcal{U}[\pi/10,\pi/2]$.
+To generate the high-fidelity (ground-truth) dataset, we sample 220 material sets, which are split into 200/10/10 for training/validation/test, respectively.
+For each material set, we sample $50$ different vertical traction conditions $T_{y}(\bm{x})$ on the top edge from a random field, following the algorithm in Lang & Potthoff (2011b); Yin et al.
+(2022).
+$T_{y}(\bm{x})$ is taken as the restriction of a 2D random field, $\phi(\bm{x})=\mathcal{F}^{-1}(\gamma^{1/2}\mathcal{F}(\Gamma))(\bm{x})$, on the top edge.
+Here, $\Gamma(\bm{x})$ is a Gaussian white noise random field on $\mathbb{R}^{2}$, $\gamma=(w_{1}^{2}+w^{2}_{2})^{-\frac{2}{4}}$ represents a correlation function, and $w_{1}$, $w_{2}$ are the wave numbers on $x$ and $y$ directions, respectively.
+Then, for each sampled traction loading, we solve the displacement field on the entire domain by minimizing the potential energy using the finite element method implemented in FEniCS (Alnæs et al., 2015).
+Sample data of the obtained dataset is illustrated in Figure 9.
+In this setting, the model takes as input the padded traction field and learns to predict the resulting displacement field.
 Figure 9: Illustration of the HGO data, loading and displacement pairs of three materials $\bm{b}^{\eta}$, $\eta=1,2,3$.
@@
 Bottom: corresponding displacement solutions $\bm{u}^{\eta}_{i}$, illustrating the impacts of system ($\bm{b}^{\eta}$) variability in solution operators.
-C.2 Example 2 - Mechanical MNIST benchmark
+C.2 Experiment 2 - Mechanical MNIST benchmark
 Mechanical MNIST is a benchmark dataset of heterogeneous material undergoing large deformation, modeled by the Neo-Hookean material with a varying modulus converted from the MNIST bitmap images (Lejeune, 2020).
@@
 (c): displacement fields (second row) $\bm{u}^{\eta}_{138}$ corresponding to the same loading field (first row) $\bm{f}_{138}$.
-C.3 Example 3 - Unsupervised heterogeneous material learning
+C.3 Experiment 3 - Unsupervised heterogeneous material learning
 We generate two sets of datasets in this case, varying the material microstructure in the following two ways.
@@
 Exemplar samples from this dataset are illustrated in Figure 11.
-These samples demonstrate the variability of $\bm{b}$, which is of critical for the latent variable identifiability in unsupervised learning settings, as proved in Theorem 2.
+These samples demonstrate the variability of $\bm{b}$, which is critical for the latent variable identifiability in unsupervised learning settings, as proved in Theorem 2.
 Varying fiber orientation magnitude and segmentation line rotation.
 Instead of controlling the fiber orientation angles on the binary field as two constant values, we sample the orientation distribution consisting of two segments with orientations $\alpha_{1}$ and $\alpha_{2}$ on each side, respectively, separated by a line passing through the center.
-The values of $\alpha_{1}$ and $\alpha_{2}$ are independently sampled from a uniform distribution over $[0,2\pi]$, and the centerline’s rotation is sampled from $[0,\pi]$.
+The values of $\alpha_{1}$ and $\alpha_{2}$ are independently sampled from a uniform distribution over $0,2\pi]$, and the centerline’s rotation is sampled from $0,\pi]$.
 We generate 300 material sets, each with 500 loading/displacement pairs, and divide these into training, validation, and test sets in a 200/50/50 split.
@@
 Here, $\Gamma(\bm{x})$ is a Gaussian white noise random field on $\mathbb{R}^{2}$, $\gamma=(w_{1}^{2}+w^{2}_{2})^{-\frac{5}{4}}$ represents a correlation function, $w_{1}$ and $w_{2}$ are the wave numbers on $x$ and $y$ directions, respectively, and $\mathcal{F}$ and $\mathcal{F}^{-1}$ denote the Fourier transform and its inverse, respectively.
-This random field is anticipated to have a zero mean and covariance operator $C=(-\Delta)^{-2.5}$, with $\Delta$ being the Laplacian with periodic boundary condition on $[0,2]^{2}$, and we then further restrict it to $\Omega$.
-For the detailed implementation of Gaussian random field sample generation, we refer interested readers to Lang and Potthoff (2011a).
-Then, for each sampled loading $\bm{f}_{i}(\bm{x})$ and microstructure field $\bm{b}^{\eta}(\bm{x})$, we solve for the displacement field $\bm{u}^{\eta}_{i}(\bm{x})$ on the entire domain.
+This random field is anticipated to have a zero mean and covariance operator $C=(-\Delta)^{-2.5}$, with $\Delta$ being the Laplacian with periodic boundary conditions on $0,2]^{2}$, and we then further restrict it to $\Omega$.
+For the detailed implementation of Gaussian random field sample generation, we refer interested readers to Lang & Potthoff (2011a).
+Then, for each sampled loading field $\bm{f}_{i}(\bm{x})$ and microstructure field $\bm{b}^{\eta}(\bm{x})$, we solve for the displacement field $\bm{u}^{\eta}_{i}(\bm{x})$ on the entire domain.
+C.4 Experiment 4 - Synthetic dataset for identifiability demonstration
+We demonstrate using experiment 4 that the disentangled latent factors from DisentangO correspond to the true generative factors.
+We generate synthetic data following the process in equation C.2.
+We work with latent variables $\bm{z}$ of dimension 2 and sample from $\bm{z}\sim\mathcal{N}(\mu_{\bm{b}},\sigma_{\bm{b}}^{2}\mathbf{I})\in\mathbb{R}^{2}$ where for each microstructure $\bm{b}$, we sample $\mu_{\bm{b}}\sim\mathcal{U}[-4,4]$ and $\sigma_{\bm{b}}\sim\mathcal{U}[1,10]$ and $f\sim\mathcal{U}[0,1]\in\mathbb{R}^{3}$, $M\sim\mathcal{U}[0,1]\in\mathbb{R}^{2\times 2}$, $A=[I,2I,3I,4I]^{T}$, $M\in\mathbb{R}^{8\times 2}$, $b\sim\mathcal{U}[0,1]\in\mathbb{R}^{8}$, $W\sim\mathcal{U}[0,1]\in\mathbb{R}^{4\times 1}$, $l\sim\mathcal{U}[0,1]\in\mathbb{R}$,
+[…]
+[…]
+We generate 700 data pairs of $(z,\theta)$ corresponding to 700 tasks, and for each task, we generate 200 samples of $(f,u)$ pairs.
+$d$ 3 5 7 9
+MCC 0.7836 0.8013 0.8237 0.9121
+$R^{2}$ 0.5673 0.6026 0.6473 0.8242
+Rank(w)(w) 3 4 4 4
+Table 5: The MCC and $R^{2}$ scores to evaluate identifiability on the synthetic dataset in experiment 4.
+Figure 12: The scatter plot of the true generative variables $z$ and the disentangled latent variables $\hat{z}$ from the synthetic dataset with $d=9$ in experiment 4.
+We conduct experiments on the synthetic dataset and split the tasks into dd/100/500 for training, validation and test, respectively, with $d$ chosen from the set of $(3,5,7,9)$ to investigate the effect of the number of training tasks on the identifiability.
+In particular, we measure the component-wise identifiability of the latent variables $\bm{z}$ by computing the Mean Correlation Coefficient (MCC) and the coefficient of determination $R^{2}$ scores on the test dataset, and report the results in Table 5, where we observe a monotonic growth on both MCC and $R^{2}$ with the increase of the number of training tasks $d$.
+Figure 12 illustrates the alignment on the test dataset between the true generative factors $z$ and the disentangled factors $\hat{z}$ from DisentangO in the case of $d=9$.
+To verify Assumption 4, we calculate the rank of the matrix $w(\bm{z},\bm{b})$.
+Specifically, for each $\theta$, with the VAE encoder, we calculate the $\mu_{\bm{b}}$ and $\sigma_{\bm{b}}$, sample $z$, and calculate the corresponding $\dfrac{\partial q_{i}(z_{i},\bm{b}^{j})}{\partial z_{i}}$ and $\dfrac{\partial^{2}q_{i}(z_{i},\bm{b}^{j})}{\partial z_{i}^{2}}$, where $i=1,\cdots d_{z},j=1\cdots d$.